### v1

# ProxyLM: Predicting Language Model Performance on Multilingual Tasks via Proxy Models

Abstract
Performance prediction is a method to estimate the performance of multilingual language models (LMs), mitigating computational costs associated with model capacity and data for fine-tuning. Our paper introduces ProxyLM, a scalable framework for predicting LM performance using proxy models in multilingual tasks. These proxy models act as surrogates, approximating the performance of fine-tuned LMs on specific downstream natural language processing (NLP) tasks. By leveraging proxy models, ProxyLM significantly reduces computational overhead on task evaluations, achieving up to a 37.08 speedup compared to traditional methods, even with our smallest proxy models. Additionally, our methodology showcases adaptability to previously unseen languages in pre-trained LMs, outperforming the state-of-the-art performance by 1.89 as measured by root-mean-square-error (RMSE). This framework streamlines model selection, enabling efficient deployment and iterative LM enhancements without extensive computational resources.
1 Introduction
Language Models (LMs) have become increasingly valuable for assessing Natural Language Processing (NLP) tasks Raffel et al. (2020); Brown et al. (2020); Costa-jussà et al. (2022b); Touvron et al. (2023a, b); Le Scao et al. (2023). However, fine-tuning and evaluating these models are resource-intensive processes in terms of both computation and time. These costs escalate with model size, especially when experimenting across multiple datasets. As highlighted in Kaplan et al. (2020), there is a scaling law that applies to both model and dataset sizes, and computational demands, indicating that larger models and broader datasets require increased computational resources. Modeling low-resource languages (LRLs) in multilingual contexts presents a range of challenges. One significant challenge is the limited data availability, which hampers effective fine-tuning processes Gu et al. (2018); Adilazuarda et al. (2024), making model adaptation through fine-tuning a challenging task Zoph et al. (2016); Liu et al. (2021). Another critical issue is the lack of pre-training data for numerous regional languages, such as Southeast Asian languages Winata et al. (2022, 2024); Yong et al. (2024), with many languages being omitted during the pre-training phase of multilingual LMs.
Given the limited academic computational resources for LM fine-tuning and inadequate LRL datasets, efficient methods in predicting model performance alleviate the dependency on extensive resources. While linear regression and gradient-boosting hold promise in performance prediction Birch et al. (2008a); Srinivasan et al. (2021); Xia et al. (2020a); Ye et al. (2021a); Schram et al. (2023); Khiu et al. (2024), existing solutions primarily focus on homogeneous data settings and prioritize high-resource languages using Transformer models Vaswani et al. (2017). Khiu et al. (2024) examine diverse datasets and LRLs but encounter limitations in the number of experiments, language diversity, and model scope, focusing solely on mBART Liu et al. (2020). Recent advancements in larger multilingual models, like NLLB Costa-jussà et al. (2022a) and M2M100 Fan et al. (2021), have significantly improved machine translation capabilities, exceeding those of mBART.
In this paper, we propose ProxyLM,111We release our code at https://github.com/davidanugraha/proxylm a framework to predict LMs performance by utilizing proxy models on LRLs. Proxy models are defined as substitute models, wherein the performance of these substitute models are used to estimate the performance of another LM. This other model can be significantly larger than our proxy models. For optimizing the prediction, we utilize much smaller LMs as proxy models and off-the-shelf models without further tuning. This approach is very scalable to multiple proxy models and task-agnostic to any modalities, thus it can be applied to any downstream tasks. This study focuses on machine translation tasks and our approach outperforms the existing work from Xia et al. (2020a); Ye et al. (2021a); Schram et al. (2023); Khiu et al. (2024), which opens a new avenue to employ LMs for model performance prediction. Therefore the contribution of our paper can be summarized in three-fold:
1.
We introduce ProxyLM, an efficient and scalable framework designed to predict the performance of LMs. This framework significantly reduces the computational costs associated with fine-tuning and inference during model selection.
2.
We demonstrate the effectiveness and robustness of ProxyLM across dataset sources and languages on two estimated LM architectures. Our framework substantially outperforms all existing baselines in English-centric, many-to-many languages, and cross-dataset settings, including scenarios involving extremely LRLs that remain unseen by pre-trained LMs, surpassing the state-of-the-art performance measured with RMSE by .
3.
We also provide a time analysis comparing the fine-tuning duration of proxy models to direct model fine-tuning. Our results indicate that, with our smallest proxy models, we can achieve up to a speedup on task evaluation compared to the traditional approach, highlighting the efficiency of our approach.
2 Methodology
In this section, we formally define the LM performance prediction problem and our proposal to improve performance prediction.
2.1 ProxyLM
Recall that performance prediction is a task of estimating a system’s performance based on the model and its training strategy, training and test dataset, and language used. Formally, let LM be our estimated model. is trained over a training dataset with source language and target language , and then tested using dataset . ’s performance, denoted , can be formulated under function that relates between these variables:
We can approximate by transforming Equation 1 into a regression task with a regressor function , which will be trained on past performance records. The regressor takes dataset features to identify the characteristics of the training and test datasets, as well as the distribution shift between them. It also takes language features to measure the dissimilarities between the source and target languages. This can be formulated as follows:
We present ProxyLM, a framework that leverages the past performance of other models, referred to as proxy models, as additional context for our regressor. Intuitively, proxy models can provide valuable insights that assist in predicting the performance of the estimated model . Formally, let be a set of proxy models. To integrate the information from these proxy models, we propose modifying Equation 2 as follows:
where represents the performance records of proxy models. The advantage of using proxy models arises from their faster fine-tuning and evaluation compared to the estimated model . This also means that off-the-shelf models can be used directly without additional tuning if they already perform the task adequately, further enhancing efficiency.
2.2 ProxyLM Features
Language Features
Dataset Features
We extract 6 features from the dataset, including train size, vocab size, average sentence length, word overlap, Type-Token Ratio (TTR), and TTR distance from and based on Xia et al. (2020a). We will refer to these features and language features combined as NLPerf features. Furthermore, we incorporate the distribution shift information between the training and test datasets using Jensen-Shannon Divergence (JSD) as described by Khiu et al. (2024). In addition, we include term frequency-inverse document frequency (TF-IDF) and sentence similarity with Sentence-BERT Reimers and Gurevych (2019).
Proxy Models Features
We leverage the performance data from proxy models, derived by averaging results from multiple fine-tuning and evaluation iterations on identical datasets and languages. Moreover, we retain the flexibility to adjust the number of proxy models employed, facilitating efficient and adaptable performance estimation.
3 Experimental Setup
In this section, we describe the datasets and LMs used to obtain LMs’ performance records. These records were then used to train various regressor models under different experimental settings to investigate our approach to performance predictions. The details of the hyper-parameters for both the LMs and the regressors are provided in A.3.
3.1 Datasets
We evaluate our approach through two machine translation benchmarks: MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023). The MT560 dataset is English-centric, where English can serve as either the source or target language. We curated 20 datasets and selected 44 languages out of 500 for evaluation in MT560. In contrast, the NusaTranslation dataset comprises parallel texts in 12 Indonesian regional languages within a Many-to-Many Languages setting, allowing any language to act as the source or target. As many of these languages are absent in pre-trained multilingual models, we analyze 8 out of the 12 languages due to limited data in the remaining 4. The datasets encompass 50 languages across various domains such as economics, technology, and medicine. Detailed language insights are available in Tables 8 and 9 in A.1 for reference.
3.2 Models
Estimated LMs
Proxy Models
We utilize four different transformer-based models: an encoder-decoder random initialized Transformers (100M) Vaswani et al. (2017), SMaLL-100 (330M) Mohammadshahi et al. (2022), M2M100 Fan et al. (2021), and NLLB Costa-jussà et al. (2022b). For M2M100 and NLLB, we use the models without any additional tuning (No FT) in a zero-shot fashion. Model details are provided in Appendix A.2. The evaluation is done using SentencePiece BLEU (spBLEU) Goyal et al. (2022), as it has been demonstrated to be a fair metric in multilingual settings, particularly in low-resource settings. For simplicity, the term “fine-tuning" will be used throughout this paper to refer to both the process of training from scratch (as in the case of the Transformer (100M) model) and the process of fine-tuning pre-trained LMs.
Regressors
We utilize XGBoost Chen and Guestrin (2016), LGBM Ye et al. (2021a), Poly2 Khiu et al. (2024), and Poly3 Khiu et al. (2024) as our regressors. In most of our experiments, we apply XGBoost as our default regressor because we find it to be the best-performing model, while the other regressors serve as baselines. Specifically for the Many-to-Many Languages setting, Matrix Factorization with context features (MF) is used as an additional baseline Schram et al. (2023). We do not apply MF to our English-centric setting because MF requires the performance records to be structured in two dimensions—one for the source language and one for the target language. In the English-centric setting, this would result in a sparse matrix with only one fully populated row or column, corresponding to English, making MF impractical for this setup.
| Models | Random | LOLO | Unseen | |||
| M2M100 | NLLB | M2M100 | NLLB | M2M100 | Avg. | |
| NLPerf Xia et al. (2020a) with different regressors | ||||||
| XGBoost | 8.09 0.44 | 8.40 0.41 | 8.81 0.39 | 12.35 0.52 | 8.22 0.46 | 9.17 |
| Poly2 Khiu et al. (2024) | 10.00 0.46 | 12.30 0.48 | 10.54 0.00 | 13.28 0.00 | 9.91 0.00 | 11.21 |
| Poly3 Khiu et al. (2024) | 9.93 0.52 | 12.33 0.46 | 10.66 0.00 | 14.06 0.00 | 10.57 0.00 | 11.51 |
| LGBM Ye et al. (2021a) | 8.18 0.34 | 8.59 0.33 | 9.00 0.49 | 12.41 0.53 | 8.59 0.69 | 9.35 |
| ProxyLM (Ours)‡ with different proxy models | ||||||
| Transformer | 5.29 0.38 | 7.65 0.47 | 6.25 0.29 | 11.33 0.50 | 6.54 0.30 | 7.41 |
| SMaLL-100 | 4.26 0.35 | 7.02 0.44 | 4.33 0.24 | 9.70 0.37 | 4.54 0.41 | 5.97 |
| SMaLL-100 (No FT) | 5.93 0.23 | 6.73 0.49 | 6.33 0.28 | 10.66 0.44 | 5.77 6.06 | 7.08 |
| Estimated Model (No FT) | 5.76 0.19 | 4.29 0.31 | 6.23 0.34 | 4.59 0.22 | 5.67 0.41 | 5.31 |
| Ensemble† | 3.79 0.30 | 3.79 0.25 | 3.90 0.22 | 4.14 0.24 | 4.36 0.22 | 4.00 |
| Models | M2M100 | NLLB | ||
| LRL | MRL | LRL | MRL | |
| Transformer | 6.14 | 6.67 | 11.60 | 10.26 |
| SMaLL-100 | 4.28 | 4.55 | 10.18 | 7.83 |
| SMaLL-100 (No FT) | 6.17 | 6.96 | 7.90 | 11.38 |
| Estimated Model (No FT) | 6.15 | 6.57 | 4.66 | 4.32 |
| Ensemble | 3.93 | 3.80 | 4.10 | 4.15 |
| Models | Random | LOLO | |||
| M2M100 | NLLB | M2M100 | NLLB | Avg. | |
| NLPerf Xia et al. (2020a) with different regressors | |||||
| XGBoost | 2.45 0.30 | 1.21 0.29 | 7.83 0.23 | 8.28 0.31 | 4.94 |
| Poly2 Khiu et al. (2024) | 4.70 0.40 | 4.68 0.51 | 7.07 0.00 | 7.90 0.00 | 6.09 |
| Poly3 Khiu et al. (2024) | 4.60 0.41 | 4.64 0.49 | 7.26 0.00 | 8.01 0.00 | 6.13 |
| LGBM Ye et al. (2021a) | 2.66 0.22 | 1.76 0.29 | 7.91 0.01 | 8.06 0.00 | 5.10 |
| MF Schram et al. (2023) | 3.65 0.26 | 2.60 0.39 | 7.08 0.23 | 7.14 0.22 | 5.12 |
| ProxyLM‡ (Ours) with different proxy models | |||||
| Transformer | 2.56 0.43 | 1.70 0.20 | 5.65 0.23 | 6.24 0.34 | 4.04 |
| SMaLL-100 | 2.56 0.33 | 1.65 0.44 | 4.85 0.36 | 5.14 0.46 | 3.55 |
| SMaLL-100 (No FT) | 2.44 0.21 | 1.34 0.38 | 6.93 0.34 | 7.25 0.37 | 4.49 |
| Estimated Model (No FT) | 2.38 0.36 | 1.27 0.30 | 5.10 0.28 | 5.50 0.26 | 3.56 |
| Ensemble† | 2.41 0.28 | 1.56 0.35 | 3.73 0.23 | 3.79 0.19 | 2.90 |
3.3 Experimental Settings
Each regressor is evaluated using RMSE as our performance metric and evaluated times with different seeds to obtain the mean and standard deviation of the performance results. We set our experiment settings as follows:
Random: We randomly sample the performance records into training and test sets with a ratio of 7:3. Then, we run 10-fold cross-validation on the training set to find the best hyper-parameters for each regressor. The best-performing regressor would subsequently be evaluated on the test set.
Leave-One-Language-Out (LOLO): We select one language as the test set, which is not encountered during training.
Unseen: The performance records can be divided into two categories: (1) records with “seen" languages and (2) records with “unseen" languages. “Unseen" languages refer to languages that are not present in the pre-training LM data, while “seen" languages denote those that are present. In this setting, the regressor is trained using records of “seen" languages and tested using records of “unseen" languages.
Cross-Dataset: We train the regressor using performance records from the MT560 dataset and test it using records from the NusaTranslation dataset. We opt not to reverse this setup as the dataset exhibits no domain shift and contains fewer performance records.
4 Results and Analysis
In this section, we present the results of the performance predictions for ProxyLM and baselines over three settings: English-centric, Many-to-Many Languages, and Cross-Dataset, as described above. Further, we discuss the robustness, effectiveness, and efficiency of ProxyLM in the context of performance prediction.
4.1 English-centric Results
Table 1 shows the overall results on MT560. ProxyLM remarkably outperforms all existing baselines. We find that incorporating all proxy models (Ensemble) is the most effective for prediction, leading to a 2.29 averaged reduction in RMSE across all experimental settings compared to the best baseline. We observe that using the “No FT" estimated model to predict the performance of their fine-tuned models is surprisingly useful in all settings, especially for NLLB, where the model already has decent machine translation quality on LRLs. This observation is supported by our findings within the XGBoost model that the NLLB No FT feature has the highest importance score among all features, as shown in Figure 16. Further, using SMaLL-100 fine-tuned performance provides useful estimations for settings involving M2M100 as the estimated model. This may indicate that the performance of a model with similar architecture can be a good estimator for the performance of the larger estimated model. In other words, the choice of proxy model to help prediction matters. Feature importance analysis from the XGBoost model supports this, revealing that the SMaLL-100 fine-tuned feature has the highest importance score among all features, as shown in Figure 15.
Our analysis also indicates that XGBoost outperforms other regression models, across all evaluated settings. Both XGBoost and LGBM, which are gradient-boosting and tree-based learning methods, demonstrate superior performance metrics across all settings. Their robustness and efficiency as non-linear models are evident when compared to linear models, such as Poly2 and Poly3. Poly2 and Poly3 regressors, which employ second-degree and third-degree polynomial regression approaches respectively, tend to generate lower scores. This diminished performance is largely attributed to their limitations in capturing the nonlinear relationships inherent in the data, leading to suboptimal results. We further present the results by the language vitality on Table 2. The overall difference in RMSE between LRLs and medium-resource languages (MRLs) is relatively small, except when using SMaLL-100 as proxy models. An interesting observation here on SMaLL-100 in predicting the NLLB model is that the “No FT" model can predict “LRL" much better than the fine-tuned counterpart, and the fine-tuned model better predicts “MRL".
| Models | M2M100 | NLLB |
| NLPerf Xia et al. (2020a) with different regressors | ||
| XGBoost | 9.97 0.35 | 10.52 0.73 |
| Poly2 Khiu et al. (2024) | 9.27 0.00 | 21.43 0.00 |
| Poly3 Khiu et al. (2024) | 8.58 0.00 | 19.81 0.00 |
| LGBM Ye et al. (2021a) | 9.24 0.24 | 15.19 1.00 |
| ProxyLM‡ (Ours) with different proxy models | ||
| Transformer | 6.65 0.26 | 10.69 0.90 |
| SMaLL-100 | 5.53 0.98 | 7.68 2.11 |
| SMaLL-100 (No FT) | 9.08 0.63 | 16.66 1.08 |
| Estimated Model (No FT) | 9.45 0.78 | 7.10 0.69 |
| Ensemble† | 4.30 0.17 | 6.21 0.50 |
4.2 Many-to-Many Languages Results
Table 3 presents the performance of different models on the NusaTranslation dataset within the Many-to-Many languages setting. The results reveal that the Ensemble model achieves the lowest RMSE, with a 1.70 averaged reduction in RMSE across all experimental settings compared to the best baseline, indicating superior accuracy in performance predictions. An exception occurs in the random NLLB setting, where the model utilizing only NLPerf features outperforms the ensemble model, achieving the best performance. Note that no domain shift occurs within the dataset.
A comparative analysis shows that predicting the performance of M2M100 models in the random setting presents a greater challenge compared to predicting for the NLLB models. This discrepancy suggests that the complexity of performance prediction can vary substantially depending on the specific LM and the conditions under which it is evaluated. A particularly noteworthy finding is the effectiveness of using “No FT" models for estimating LM performance. The “No FT" models, which do not require any additional fine-tuning, demonstrate high accuracy in their performance predictions. This method offers substantial efficiency benefits, as it eliminates the need for extensive computational resources typically required for model training. In contrast, we find similar results between the LOLO setting for Many-to-Many languages and English-centric results, where ProxyLM using Ensemble remarkably outperforms all existing baselines. In addition, we find that using SMaLL-100 fine-tuned performance results in better predictions compared to those of the “No FT" estimated model.
4.3 Cross-Dataset Results
Table 4 illustrates model performance in the Cross-Dataset setup, showcasing the superior performance of ProxyLM with LGBM over XGBoost. The results highlight that ProxyLM with Ensemble significantly reduces RMSE compared to the best baseline by 2 and 1.69 for M2M100 and NLLB, respectively. This displays consistent performance across datasets and languages that were not encountered during the regressor’s training, including “unseen" languages for the pre-trained LMs. Moreover, the “No FT" models exhibit variability compared to other proxy models. The performance variation between M2M100 and NLLB may be attributed to the MT560 dataset solely containing "seen" languages for NLLB, lacking “unseen" languages examples for the regressor. This highlights the significance of incorporating “unseen" language instances in training for more dependable predictions.
4.4 Ablation Study
Figure 2 highlights the impact of features used in ProxyLM in the LOLO setting with XGBoost. Utilizing proxy models as features leads to a significant reduction in RMSE across all scenarios, showcasing their importance compared to other features. For the MT560 dataset, including language and dataset features alongside proxy models enhances performance. Dataset features alone show better improvement than language features alone, but the combination of both yields the best performance. On the other hand, for the NusaTranslation dataset, the benefits of incorporating dataset and language features are less pronounced, especially for the M2M100 model, and there may even be a performance dip for the NLLB model due to the dataset’s lack of domain shift.
4.5 Diminishing Returns with Increasing Training Set Size
In Figure 3, we examine the training of the XGBoost regressor using different numbers of MT560 past performance records as the training dataset. While the regressor’s performance shows enhancement with an expanding training size, the incremental benefits start diminishing once the training set surpasses about 400 past performance records. This observation implies that, across datasets, there exists a threshold where the advantages of incorporating additional past performance records begin to exhibit diminishing returns.
| Datasets | Inference | Fine-tuning | |||
| MT560 | NusaTranslation | ||||
| Estimated models | |||||
| M2M100 | 338 s | 3.94 | hrs (4.77) | 1.42 | hrs (7.32) |
| NLLB | 492 s | 9.4 | hrs (11.48) | 7.21 | hrs (37.08) |
| Proxy models | |||||
| SMaLL-100 | 201 s | 2.5 | hrs (3.03) | 1.03 | hrs (5.29) |
| Transformer | 60 s | 0.83 | hrs (1) | 0.19 | hrs (1) |
| Regressor Model | MT560 | NusaTranslation |
| XGBoost | 129.29 s | 69.81 s |
| Poly2 | 0.06 s | 0.03 s |
| Poly3 | 0.07 s | 0.06 s |
| LGBM | 702.12 s | 173.17 s |
| MF | N/A | 465.37 s |
4.6 Time Efficiency
Table 5 compares the fine-tuning and inference times required for the estimated and proxy models. The results demonstrate that fine-tuning proxy models or direct inference from any model is remarkably faster than fine-tuning all estimated models. Table 6 further illustrates this point, showing only a minimal trade-off in the time needed to train the regressor models. This additional training time is relatively negligible, highlighting the efficiency of using proxy models.
4.7 Performance by Language Categories
In Figure 4, we present detailed XGBoost results with ProxyLM Ensemble on the M2M100 model under the English-centric LOLO experiment, grouped by language categories. Based on the Locally Weighted Scatterplot Smoothing (LOWESS) Cleveland (1979) curve depicted in Figure 4(c), our method consistently maintains unbiased predictions for spBLEU scores below 40 across various language types. However, as the spBLEU score increases, the availability of data points diminishes, leading to our method under-predicting the performance compared to the true spBLEU score. Outliers observed in Kartvelian languages and Indo-European languages with Joshi class 3 may have contributed to this discrepancy in prediction. These observations suggest that increasing the number of data points covering higher spBLEU scores may help mitigate the bias in prediction. Further experiment details are available in Appendix B.
5 Related Work
The prediction performance of machine learning algorithms has been mainly explored in two research directions: (1) predict the model performance during the training runtime, and (2) predict the model performance by providing extracted features from the dataset Xia et al. (2020a).
Performance Prediction During the Training Runtime
The former aims to infer and extrapolate the learning curve to approximate training results using evaluation metric measurements Kolachina et al. (2012). Domhan et al. (2015) study the quick detection of poor hyper-parameters in probabilistic models after a few steps of Stochastic Gradient Descent (SGD). Adriaensen et al. (2024) extrapolate learning curves from a parametric prior using Markov Chain Monte Carlo (MCMC).
Performance Prediction Using Extracted Features
The latter aims to predict the model performance by learning a correlation between input features and final evaluation metric. Birch et al. (2008a) identify strong predictive features such as the amount of reordering, the morphological complexity of the target language, and the historical relatedness of the two languages. Xia et al. (2020a) leverage extracted dataset features and typological database language representations. Ye et al. (2021a) introduce the use of confidence intervals and calibration with various regressor algorithms for reliable performance prediction. Schram et al. (2023) apply Bayesian matrix factorization for performance prediction on multilingual NLP tasks. In this work, we focus to explore the latter. Existing approaches have shown promise using linear regression and gradient-boosting trees Birch et al. (2008b); Xia et al. (2020b); Srinivasan et al. (2021); Ye et al. (2021b). These studies have considered data size, typological features, and language similarity as factors contributing to the model performance.
6 Conclusion
In this paper, we introduce ProxyLM, a novel framework designed to predict the performance of LMs by leveraging proxy models specifically for LRLs. By utilizing proxy models as substitutes to estimate the performance of the target model, we strategically employ smaller LMs and off-the-shelf models without additional fine-tuning. This framework is highly scalable to multiple proxy models and is task-agnostic, making it applicable to a wide range of downstream tasks. Our streamlined approach showcases substantial advancements in prediction accuracy compared to standard baselines and exhibits strong generalization capabilities across varied scenarios.
Limitations
This paper focuses exclusively on two estimated models: M2M100 and NLLB, to evaluate our proposed framework. For demonstration purposes, we concentrate on the usage of specific models, namely the Transformer model, the SMaLL-100 model, and the No FT models, to illustrate the effectiveness of our proxy models. The M2M100 and NLLB models were selected due to their prominence and relevance in the field of multilingual translation tasks. These models serve as robust benchmarks for assessing the performance and reliability of our proxy-based framework. By using these well-regarded models, we aim to provide compelling evidence of the capabilities and advantages of ProxyLM. While our proposed framework is evaluated solely within the context of machine translation, it is not confined to this application alone. The framework is designed to be versatile and can be extended to a variety of other downstream tasks. We plan to explore these additional applications in future work.
Some other possible avenues for future work could involve a deeper investigation into which proxy models are more effective for enhancing performance prediction in specific settings. Our findings suggest that one proxy model can outperform another in different scenarios, making it crucial to carefully select the most relevant proxy models to maximize the benefits of our approach. Additionally, developing methodologies for collecting relevant past performance records could provide better insights and improve the generalization and accuracy of our framework. Past performance records may provide better information gain than others, potentially minimizing the number of performance records required for a more robust and accurate predictor.
Acknowledgements
We extend our sincere gratitude to Viktoria Schram for providing assistance to reproduce baselines.
Appendix A Experimental Details
| Datasets | Languages Under Study | Domain |
| MT560 Gowda et al. (2021) | ||
| Joshua Indian Corpus | tam | Wiki |
| Neulab TED Talk Qi et al. (2018) | bel, hye, kat, kaz, mar, mya, tam | TED |
| News Commentary | kaz | Economics |
| OPUS100 Zhang et al. (2020) | afr, amh, bel, cym, guj, hau, hye, ibo, ind, kan, | Multi-domain |
| kat, kaz, khm, kin, kir, mar, mya, oci, pan, sin, | ||
| tam, tat, tuk, xho, yor, zul | ||
| OPUS Bible | afr, amh, ceb, dik, ewe, guj, hye, ind, kan, mar, | Religion |
| Christodouloupoulos and Steedman (2015) | mri, mya, plt, sin, sna, som, ssw, tam, tgl, wol, | |
| xho, zul | ||
| OPUS OpenSubtitles | afr, hye, kat, sin, tgl | Movies |
| OPUS Tatoeba Tiedemann (2012) | afr, amh, arz, bak, bel, ceb, fao, hau, hye, ibo, | Conversational |
| ind, jav, kan, kat, kaz, khm, kin, kir, lmo, ltz, | ||
| mar, mri, mya, pan, sna, tat, tgl, tuk, yor, zul | ||
| OPUS Tanzil Tiedemann (2012) | amh, hau, ind, snd, som, tat | Religion |
| OPUS Gnome Tiedemann (2012) | fao, mri, som | Technology |
| OPUS GlobalVoices Tiedemann (2012) | amh, mya | News |
| OPUS EU Bookshop Tiedemann (2012) | cym | Government |
| OPUS SPC | afr | Government |
| OPUS Memat | xho | Medicine |
| OPUS XhosaNavy | xho | Government |
| OPUS KDE4 | hne | Technology |
| OPUS infopankki | som | Immigration |
| PMIndia Haddow and Kirefu (2020) | guj, kan, mar, pan, sin, tam | Government |
| Wiki Titles | guj, kaz, tam | Wiki |
| WikiMatrix Schwenk et al. (2021) | arz, bak, bel, ceb, fao, ind, jav, kat, kaz, lmo, | Wiki |
| ltz, mar, oci, sin, tam, tat, tgl | ||
| FLoRes Costa-jussà et al. (2022b) | afr, amh, arz, bak, bel, ceb, cym, dik, ewe, fao, | Multi-domain |
| guj, hau, hne, hye, ibo, ind, jav, kan, kat, kaz, | ||
| khm, kin, kir, lmo, ltz, mar, mri, mya, oci, pan, | ||
| plt, sin, sna, snd, som, ssw, tam, tat, tgl, tuk, | ||
| wol, xho, yor, zul | ||
| NusaTranslation Cahyawijaya et al. (2023) | bew, btk, ind, jav, mad, mak, min, sun | Social Media |
A.1 Languages Under Study
We list all the languages used in the training from the MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023) datasets in Table 8 and Table 9, respectively. The language code follows ∗ISO639-3 coding. All languages are also complemented by their †rarity taxonomy based on Joshi et al. (2021) into two vitality classes: 0-2→low resource language (LRL), and 3→mid resource language (MRL). We also provide information about whether the language was part of the pretrained M2M100 model dataset to highlight the model knowledge coverage.
A.2 Models
Here are the details on the proxy LMs we use in the experiments as follows:
Transformer (100M) Vaswani et al. (2017): a standard encoder-decoder transformer-based model with 6 encoder layers and 6 decoder layers with an embedding dimension of 512. We train the model from randomly initialized parameters with the training set.
SMaLL-100 (330M) Mohammadshahi et al. (2022):222SMaLL-100 (330M) is taken from https://github.com/alirezamshi/small100. a distilled version of the M2M100 (12B) model. We utilize the model in two ways: fine-tuned on training data and zero-shot inference.
M2M100 (No FT) Fan et al. (2021):333M2M100 (1.2B) is taken from https://github.com/facebookresearch/fairseq/tree/main/examples/m2m_100. a pre-trained estimated model of M2M100 (1.2B) without any fine-tuning. We run the model in a zero-shot fashion.
NLLB (No FT) Costa-jussà et al. (2022b):444NLLB (1.3B) is taken from https://github.com/facebookresearch/fairseq/tree/nllb. a pre-trained estimated model of NLLB-200 Distilled (1.3B) without any fine-tuning. We run the model in a zero-shot fashion.
| Language | Language Code∗ | Family | Joshi Class† | Vitality† | Seen by M2M100 |
| Afrikaans | afr | indo-european | 3 | MRL | ✓ |
| Amharic | amh | afro-asiatic | 2 | LRL | ✓ |
| Armenian | hye | indo-european | 1 | LRL | ✓ |
| Bashkir | bak | turkic | 1 | LRL | ✓ |
| Belarusian | bel | indo-european | 3 | MRL | ✓ |
| Burmese | mya | sino-tibetan | 1 | LRL | ✓ |
| Cebuano | ceb | austronesian | 3 | MRL | ✓ |
| Chhattisgarhi | hne | indo-european | 0 | LRL | ✗ |
| Dinka | dik | nilo-saharan | 1 | LRL | ✗ |
| Egyptian Arabic | arz | afro-asiatic | 3 | MRL | ✓ |
| Ewe | ewe | niger-congo | 1 | LRL | ✗ |
| Faroese | fao | indo-european | 1 | LRL | ✗ |
| Georgian | kat | kartvelian | 3 | MRL | ✓ |
| Gujarati | guj | indo-european | 1 | LRL | ✓ |
| Hausa | hau | afro-asiatic | 2 | LRL | ✓ |
| Igbo | ibo | niger-congo | 1 | LRL | ✓ |
| Indonesian | ind | austronesian | 3 | MRL | ✓ |
| Javanese | jav | austronesian | 1 | LRL | ✓ |
| Kannada | kan | dravidian | 1 | LRL | ✓ |
| Kazakh | kaz | turkic | 3 | MRL | ✓ |
| Khmer | khm | austro-asiatic | 1 | LRL | ✓ |
| Kirghiz | kir | turkic | 1 | LRL | ✗ |
| Kinyarwanda | kin | niger-congo | 1 | LRL | ✗ |
| Lombard | lmo | indo-european | 1 | LRL | ✗ |
| Luxembourgish | ltz | indo-european | 1 | LRL | ✓ |
| Malagasy | plt | austronesian | 1 | LRL | ✓ |
| Maori | mri | austronesian | 1 | LRL | ✗ |
| Marathi | mar | indo-european | 2 | LRL | ✓ |
| Occitan | oci | indo-european | 1 | LRL | ✓ |
| Punjabi | pan | indo-european | 2 | LRL | ✓ |
| Shona | sna | niger-congo | 1 | LRL | ✗ |
| Sindhi | snd | indo-european | 1 | LRL | ✓ |
| Sinhala | sin | indo-european | 0 | LRL | ✓ |
| Somali | som | afro-asiatic | 1 | LRL | ✓ |
| Swati | ssw | niger-congo | 1 | LRL | ✓ |
| Tagalog | tgl | austronesian | 3 | MRL | ✓ |
| Tamil | tam | dravidian | 3 | MRL | ✓ |
| Tatar | tat | turkic | 1 | LRL | ✗ |
| Turkmen | tuk | turkic | 1 | LRL | ✗ |
| Welsh | cym | indo-european | 1 | LRL | ✓ |
| Wolof | wol | niger-congo | 2 | LRL | ✓ |
| Xhosa | xho | niger-congo | 2 | LRL | ✓ |
| Yoruba | yor | niger-congo | 2 | LRL | ✓ |
| Zulu | zul | niger-congo | 2 | LRL | ✓ |
A.3 Hyper-parameters
LM
Each fine-tuning and evaluation for LMs is done with an NVIDIA Tesla V100 32GB GPU. The hyper-parameters used during fine-tuning from the MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023) datasets are listed in Table 10, 11, 12, and 13 for SMaLL100, M2M100, NLLB, and Transformer models, respectively.
Regressor
Each regressor is trained on an AMD Ryzen Threadripper 2990WX with 128 GB of RAM and 16 threads. Regressors’ hyper-parameters used are provided in Table 14, 15, 16, and 17 for XGB, Poly2/Poly3, LGBM, and MF, respectively. These hyper-parameters were obtained based on the best cross-validation RMSE score using 10 folds.
| Language | Language Code∗ | Family | Joshi Class† | Vitality† | Seen by M2M100 | Seen by NLLB |
| Indonesian | ind | austronesian | 3 | MRL | ✓ | ✓ |
| Javanese | jav | austronesian | 1 | LRL | ✓ | ✓ |
| Betawi | bew | creole | 0 | LRL | ✗ | ✗ |
| Batak | btk | austronesian | 0 | LRL | ✗ | ✗ |
| Madurese | mad | austronesian | 0 | LRL | ✗ | ✗ |
| Makassarese | mak | austronesian | 0 | LRL | ✗ | ✗ |
| Minangkabau | min | austronesian | 0 | LRL | ✗ | ✓ |
| Sundanese | sun | austronesian | 1 | LRL | ✓ | ✓ |
Appendix B More Detailed Results
We provide detailed visualizations of the results of XGBoost with ProxyLM Ensemble based on multiple language groupings in Figure 5 - 10 for English-centric result, and Figure 11 - 14 for Many-to-Many Languages result. Each language groupings plot comprises multiple subplots, including (a) vitality class, (b) Joshi class, (c) language family, and (d) individual languages. The mapping of vitality, Joshi class, and language family follows the classifications in Table 8 and 9.
| Hyper-parameter | MT560 | NusaTranslation |
| Encoder Layers | 12 | 12 |
| Decoder Layers | 3 | 3 |
| Encoder Embed Dim | 1024 | 1024 |
| Decoder Embed Dim | 1024 | 1024 |
| Encoder FFN Embed Dim | 4096 | 4096 |
| Decoder FFN Embed Dim | 4096 | 4096 |
| Encoder Attention Heads | 16 | 16 |
| Decoder Attention Heads | 16 | 16 |
| Encoder Layerdrop | 0.05 | 0.05 |
| Decoder Layerdrop | 0.05 | 0.05 |
| Optimizer | Adam | Adam |
| Adam Eps | 1e-6 | 1e-6 |
| Adam Betas | (0.9, 0.98) | (0.9, 0.98) |
| Patience | 6 | 6 |
| Batch Size | 16 | 16 |
| Dropout | 0.1 | 0.1 |
| Attention Dropout | 0.1 | 0.1 |
| ReLU Dropout | 0.1 | 0.1 |
| Weight Decay | 0.0 | 0.0 |
| Label Smoothing | 0.1 | 0.1 |
| Clip Norm | 1.0 | 1.0 |
| Learning Rate | 0.0001 | 0.0003 |
| Max Tokens (per GPU) | 1,000 | 1,000 |
| Hyper-parameter | MT560 | NusaTranslation |
| Encoder Layers | 24 | 24 |
| Decoder Layers | 24 | 24 |
| Encoder Embed Dim | 1024 | 1024 |
| Decoder Embed Dim | 1024 | 1024 |
| Encoder FFN Embed Dim | 8192 | 8192 |
| Decoder FFN Embed Dim | 8192 | 8192 |
| Encoder Attention Heads | 16 | 16 |
| Decoder Attention Heads | 16 | 16 |
| Encoder Layerdrop | 0.05 | 0.05 |
| Decoder Layerdrop | 0.05 | 0.05 |
| Optimizer | Adam | Adam |
| Adam Eps | 1e-6 | 1e-6 |
| Adam Betas | (0.9, 0.98) | (0.9, 0.98) |
| Patience | 6 | 6 |
| Batch Size | 32 | 32 |
| Dropout | 0.1 | 0.1 |
| Attention Dropout | 0.1 | 0.1 |
| ReLU Dropout | 0.1 | 0.1 |
| Weight Decay | 0.0 | 0.0 |
| Label Smoothing | 0.1 | 0.1 |
| Clip Norm | 0.0 | 0.0 |
| Learning Rate | 0.0002 | 0.0002 |
| Max Tokens (per GPU) | 1,792 | 1,792 |
Appendix C Further Analysis
C.1 Model Feature Importance
We provide feature importance scores of XGBoost with ProxyLM Ensemble for the random MT560 experiment in Figure 15 and 16. Each combination consists of one most influential feature followed by others with marginal contributions to the model, each with an importance score of 0.12 or less. We observe that proxy models are always the most influential features in prediction.
| Hyper-parameter | MT560 | NusaTranslation |
| Encoder Layers | 24 | 24 |
| Decoder Layers | 24 | 24 |
| Encoder Embed Dim | 1024 | 1024 |
| Decoder Embed Dim | 1024 | 1024 |
| Encoder FFN Embed Dim | 8192 | 8192 |
| Decoder FFN Embed Dim | 8192 | 8192 |
| Encoder Attention Heads | 16 | 16 |
| Decoder Attention Heads | 16 | 16 |
| Encoder Layerdrop | 0.05 | 0.05 |
| Decoder Layerdrop | 0.05 | 0.05 |
| Optimizer | Adam | Adam |
| Adam Eps | 1e-6 | 1e-6 |
| Adam Betas | (0.9, 0.98) | (0.9, 0.98) |
| Patience | 6 | 6 |
| Batch Size | 32 | 32 |
| Dropout | 0.1 | 0.1 |
| Attention Dropout | 0.1 | 0.1 |
| ReLU Dropout | 0.0 | 0.0 |
| Weight Decay | 0.01 | 0.01 |
| Label Smoothing | 0.1 | 0.1 |
| Clip Norm | 1.0 | 1.0 |
| Learning Rate | 0.00002 | 0.0001 |
| Max Tokens (per GPU) | 1,000 | 1,000 |
| Hyper-parameter | MT560 | NusaTranslation |
| Encoder Layers | 6 | 6 |
| Decoder Layers | 6 | 6 |
| Encoder Embed Dim | 512 | 512 |
| Decoder Embed Dim | 512 | 512 |
| Encoder FFN Embed Dim | 2048 | 2048 |
| Decoder FFN Embed Dim | 2048 | 2048 |
| Encoder Attention Heads | 8 | 8 |
| Decoder Attention Heads | 8 | 8 |
| Encoder Layerdrop | 0.05 | 0.05 |
| Decoder Layerdrop | 0.05 | 0.05 |
| Optimizer | Adam | Adam |
| Adam Eps | 1e-6 | 1e-6 |
| Adam Betas | (0.9, 0.98) | (0.9, 0.98) |
| Patience | 6 | 6 |
| Batch Size | 32 | 32 |
| Dropout | 0.1 | 0.1 |
| Attention Dropout | 0.1 | 0.1 |
| ReLU Dropout | 0.1 | 0.1 |
| Weight Decay | 0.0001 | 0.0001 |
| Label Smoothing | 0.1 | 0.1 |
| Clip Norm | 0 | 0 |
| Learning Rate | 0.001 | 0.0005 |
| Max Tokens (per GPU) | 1,000 | 1,000 |
| Hyper-parameter | NusaTranslation | MT560 | ||
| M2M100 | NLLB | M2M100 | NLLB | |
| max n_estimators | 2000 | 2000 | 5000 | 5000 |
| max eta | 0.1 | 0.1 | 0.1 | 0.1 |
| min_child_weight | 5 | 2.5 | 3 | 3.5 |
| max_depth | 3 | 3 | 5 | 4 |
| gamma | 0 | 0 | 0.2 | 0.1 |
| subsample | 0.7 | 0.9 | 0.95 | 0.8 |
| colsample_bytree | 0.6 | 0.6 | 0.85 | 1.0 |
| reg_alpha | 0 | 0 | 0.1 | 0.1 |
| reg_lambda | 0.35 | 0.15 | 0.1 | 0.3 |
| Hyper-parameter | Value |
| alpha | 0.1 |
| l1_ratio | 0.9 |
| Hyper-parameter | Value |
| max learning_rate | 0.3 |
| max num_leaves | 64 |
| n_estimators | 100 |
| max_bin | 200000 |
| max_depth | 10 |
| min_child_weight | 0.001 |
| min_child_samples | 20 |
| min_split_gain | 0.0 |
| colsample_bytree | 1.0 |
| subsample | 1.0 |
| reg_alpha | 0.1 |
| reg_lambda | 0.1 |
| Hyper-parameter | Specification |
| max alpha | 0.01 |
| beta_w | 0.1 |
| beta_h | 0.1 |
| beta_z | 0.01 |
| beta_s | 0.01 |
| beta_t | 0.01 |
| lr_decay | 0.001 |
| iterations | 2000 |