### v2

# ProxyLM: Predicting Language Model Performance on Multilingual Tasks via Proxy Models

Abstract
Performance prediction is a method to estimate the performance of Language Models (LMs) on various Natural Language Processing (NLP) tasks, mitigating computational costs associated with model capacity and data for fine-tuning. Our paper introduces ProxyLM, a scalable framework for predicting LM performance using proxy models in multilingual tasks. These proxy models act as surrogates, approximating the performance of the LM of interest. By leveraging proxy models, ProxyLM significantly reduces computational overhead on task evaluations, achieving up to a 37.08$\times$ speedup compared to traditional methods, even with our smallest proxy models. Additionally, our methodology showcases adaptability to previously unseen languages in pre-trained LMs, outperforming the state-of-the-art performance by 1.89$\times$ as measured by root-mean-square error (RMSE). This framework streamlines model selection, enabling efficient deployment and iterative LM enhancements without extensive computational resources.
1 Introduction
Language Models (LMs) have become increasingly valuable for assessing Natural Language Processing (NLP) tasks Raffel et al. (2020); Brown et al. (2020); Costa-jussà et al. (2022b); Touvron et al. (2023a, b); Le Scao et al. (2023). However, fine-tuning and evaluating these models are resource-intensive processes in terms of both computation and time. These costs escalate with model size, especially when experimenting across multiple datasets. As highlighted in Kaplan et al. (2020), there is a scaling law that applies to both model and dataset sizes, and computational demands, indicating that larger models and broader datasets require increased computational resources. Modeling low-resource languages (LRLs) in multilingual contexts presents a range of challenges. One significant challenge is the limited data availability, which hampers effective fine-tuning processes Gu et al. (2018); Adilazuarda et al. (2024), making model adaptation through fine-tuning a challenging task Zoph et al. (2016); Liu et al. (2021). Another critical issue is the lack of pre-training data for numerous regional languages, such as Southeast Asian languages Winata et al. (2022, 2024); Yong et al. (2024), with many languages being omitted during the pre-training phase of multilingual LMs.
Figure 1: ProxyLM framework for LM performance prediction. (Top) The evaluation metric is computed on the test set using a proxy model $\mathcal{M}_{p}^{i}$. (Bottom) The regressor $g$ is trained using proxy model scores as well as dataset and language features by minimizing the RMSE difference of $y_{\mathcal{M}}$ and $\hat{y}_{\mathcal{M}}$.
Given the limited academic computational resources for LM fine-tuning and inadequate datasets for LRLs, performance prediction is an efficient method to alleviate the dependency on extensive resources by leveraging past performance records on an NLP task. While linear regression and gradient-boosting hold promise in performance prediction Birch et al. (2008a); Srinivasan et al. (2021); Xia et al. (2020a); Ye et al. (2021a); Schram et al. (2023); Khiu et al. (2024), existing solutions primarily focus on homogeneous data settings and prioritize high-resource languages using Transformer models Vaswani et al. (2017). Khiu et al. (2024) examine diverse datasets and LRLs but encounter limitations in the number of experiments, language diversity, and model scope, focusing solely on mBART Liu et al. (2020). Recent advancements in larger multilingual models, like NLLB Costa-jussà et al. (2022a) and M2M100 Fan et al. (2021), have significantly improved machine translation capabilities, exceeding those of mBART.
In this paper, we propose ProxyLM,  [^1] We release our code at https://github.com/davidanugraha/proxylm a framework to predict LMs performance by utilizing proxy models on LRLs. Proxy models are defined as substitute models, wherein the performance of these substitute models are used to estimate the performance of another LM. This other model can be significantly larger than our proxy models. For optimizing the prediction, we utilize much smaller LMs as proxy models and off-the-shelf models without further tuning. This approach is very scalable to multiple proxy models and task-agnostic to any modalities, thus it can be applied to any downstream tasks. This study focuses on machine translation tasks and our approach outperforms the existing work from Xia et al. (2020a); Ye et al. (2021a); Schram et al. (2023); Khiu et al. (2024), which opens a new avenue to employ LMs for model performance prediction. Therefore the contribution of our paper can be summarized in three-fold:
  1. 1.
    We introduce ProxyLM, an efficient and scalable framework designed to predict the performance of LMs. This framework significantly reduces the computational costs associated with fine-tuning and inference during model selection.
  2. 2.
    We demonstrate the effectiveness and robustness of ProxyLM across 21 dataset sources and 50 languages on two estimated LM architectures. Our framework substantially outperforms all existing baselines in English-centric, many-to-many languages, and cross-dataset settings, including scenarios involving extremely LRLs that remain unseen by pre-trained LMs, surpassing the state-of-the-art performance measured with root-mean-square error (RMSE) by 1.89$\times$.
  3. 3.
    We also provide a time analysis comparing the fine-tuning duration of proxy models to direct model fine-tuning. Our results indicate that, with our smallest proxy models, we can achieve up to a 37.08$\times$ speedup on task evaluation compared to the traditional approach, highlighting the efficiency of our approach.
2 Methodology
In this section, we formally define the LM performance prediction problem and our proposal to improve performance prediction.
2.1 ProxyLM
Recall that performance prediction is a task of estimating a system’s performance based on the model and its training strategy, training and test dataset, and language used. Formally, let LM $\mathcal{M}$ be our estimated model. $\mathcal{M}$ is trained over a training dataset $\mathcal{D}$ with source language $\mathcal{L}_{s}$ and target language $\mathcal{L}_{t}$, and then tested using dataset $\mathcal{D}^{\prime}$. $\mathcal{M}$’s performance, denoted $y_{\mathcal{M}}$, can be formulated under function $f$ that relates between these variables:
$y_{\mathcal{M}}=f(\mathcal{M},\mathcal{D},\mathcal{D}^{\prime},\mathcal{L}_{s},\mathcal{L}_{t}).$ (1)
We can approximate $f$ by transforming Equation 1 into a regression task with a regressor function $g$, which will be trained on past performance records. The regressor takes dataset features $\Phi(\mathcal{D},\mathcal{D}^{\prime})$ to identify the characteristics of the training and test datasets, as well as the distribution shift between them. It also takes language features $\Psi(\mathcal{L}_{s},\mathcal{L}_{t})$ to measure the dissimilarities between the source and target languages. This can be formulated as follows:
$\hat{y_{\mathcal{M}}}=g(\Phi(\mathcal{D},\mathcal{D}^{\prime});\Psi(\mathcal{L}_{s},\mathcal{L}_{t})).$ (2)
We present ProxyLM, a framework that leverages the past performance of other models, referred to as proxy models, as additional context for our regressor. Intuitively, proxy models can provide valuable insights that assist in predicting the performance of the estimated model $\mathcal{M}$. Formally, let $\mathcal{M}_{p}=[\mathcal{M}_{p}^{1},\dots,\mathcal{M}_{p}^{N}]$ be a set of $N$ proxy models. To integrate the information from these proxy models, we modify Equation 2 as follows:
$\hat{y_{\mathcal{M}}}=g(\hat{y}_{\mathcal{M}_{p}};\Phi(\mathcal{D},\mathcal{D}^{\prime});\Psi(\mathcal{L}_{s},\mathcal{L}_{t})),$ (3)
where $y_{\mathcal{M}_{p}}=[y_{\mathcal{M}_{p}^{1}},\dots,y_{\mathcal{M}_{p}^{N}}]$ represents the performance records of $N$ proxy models. The advantage of using proxy models arises from their faster fine-tuning and evaluation compared to the estimated model $\mathcal{M}$. This also means that off-the-shelf models can be used directly without additional tuning if they already perform the task adequately, further enhancing efficiency.
2.2 ProxyLM Features
Language Features
We use URIEL Typological Database Littell et al. (2017) similar to Xia et al. (2020a) including geographic, genetic, inventory, syntactic, phonological, and featural distance. The language features are useful to provide a language-specific representation to the regressor.
Dataset Features
We extract 6 features from the dataset, including train size, vocab size, average sentence length, word overlap, Type-Token Ratio (TTR), and TTR distance from $\mathcal{D}$ and $\mathcal{D^{\prime}}$ based on  Xia et al. (2020a). We will refer to these features and language features combined as NLPerf features. Furthermore, we incorporate the distribution shift information between the training and test datasets using Jensen-Shannon Divergence (JSD) as described by  Khiu et al. (2024). In addition, we include term frequency-inverse document frequency (TF-IDF) and sentence similarity with Sentence-BERT  Reimers and Gurevych (2019).
Proxy Models Features
We leverage the performance data from proxy models, derived by averaging results from multiple fine-tuning and evaluation iterations on identical datasets and languages. Moreover, we retain the flexibility to adjust the number of proxy models employed, facilitating efficient and adaptable performance estimation.
3 Experimental Setup
In this section, we describe the datasets and LMs used to obtain LMs’ performance records. These records are then used to train various regressor models under different experimental settings to investigate our approach to performance predictions. The details of the hyper-parameters for both the LMs and the regressors are provided in A.3.
3.1 Datasets
We evaluate our approach using two types of datasets: English-centric and Many-to-Many Languages. The English-centric dataset involves English serving as either the source or target language. Our English-centric dataset comes from the MT560 Gowda et al. (2021) dataset, where we curated 19 datasets and selected 44 languages out of 500 for evaluation. Furthermore, we use the FLoRes-200 dataset Costa-jussà et al. (2022b) for additional validation and test sets. In contrast, the Many-to-Many Languages dataset allows any language to act as the source or target. We use the NusaTranslation dataset  Cahyawijaya et al. (2023) as our Many-to-Many Languages dataset, which comprises parallel texts in 12 Indonesian regional languages. As many of these languages are absent in pre-trained multilingual models, we analyze 8 out of the 12 languages due to limited data in the remaining 4. Both datasets encompass 50 languages across various domains such as economics, technology, and medicine. Detailed language insights are available in Section A.1 in the Appendix.
3.2 Models
Estimated LMs
We employ two estimated LMs: M2M100 1.2B Fan et al. (2021) and NLLB 1.3B Costa-jussà et al. (2022b). Each estimated model is fine-tuned using a standard next token prediction objective on the training set.
Proxy Models
We utilize four different transformer-based models: an encoder-decoder random initialized Transformers (100M) Vaswani et al. (2017), SMaLL-100 (330M) Mohammadshahi et al. (2022), M2M100 Fan et al. (2021), and NLLB  Costa-jussà et al. (2022b). For M2M100 and NLLB, we use the models without any additional tuning (No FT) in a zero-shot fashion. Model details are provided in Section  A.2 in the Appendix. The evaluation is done using SentencePiece BLEU (spBLEU) Goyal et al. (2022), as it has been demonstrated to be a fair metric in multilingual settings, particularly in low-resource settings. For simplicity, the term “fine-tuning" will be used throughout this paper to refer to both the process of training from scratch (as in the case of the Transformer (100M) model) and the process of fine-tuning pre-trained LMs.
Regressors
We utilize XGBoost Chen and Guestrin (2016), LGBM Ye et al. (2021a), Poly2 Khiu et al. (2024), and Poly3 Khiu et al. (2024) as our regressors. In most of our experiments, we apply XGBoost as our default regressor because we find it to be the best-performing model, while the other regressors serve as baselines. Specifically for the Many-to-Many Languages setting, Matrix Factorization with context features (MF) is used as an additional baseline Schram et al. (2023). We do not apply MF to our English-centric setting because MF requires the performance records to be structured in two dimensions—one for the source language and one for the target language. In the English-centric setting, this would result in a sparse matrix with only one fully populated row or column, corresponding to English, making MF impractical for this setup.
Models Random LOLO Unseen
[…]
NLPerf Xia et al. (2020a) with different regressors
XGBoost 8.09 $\pm$ 0.44 8.40 $\pm$ 0.41 8.81 $\pm$ 0.39 12.35 $\pm$ 0.52 8.22 $\pm$ 0.46 9.17
Poly2 Khiu et al. (2024) 10.00 $\pm$ 0.46 12.30 $\pm$ 0.48 10.54 $\pm$ 0.00 13.28 $\pm$ 0.00 9.91 $\pm$ 0.00 11.21
Poly3 Khiu et al. (2024) 9.93 $\pm$ 0.52 12.33 $\pm$ 0.46 10.66 $\pm$ 0.00 14.06 $\pm$ 0.00 10.57 $\pm$ 0.00 11.51
LGBM Ye et al. (2021a) 8.18 $\pm$ 0.34 8.59 $\pm$ 0.33 9.00 $\pm$ 0.49 12.41 $\pm$ 0.53 8.59 $\pm$ 0.69 9.35
ProxyLM (Ours) with different proxy models
Transformer 5.29 $\pm$ 0.38 7.65 $\pm$ 0.47 6.25 $\pm$ 0.29 11.33 $\pm$ 0.50 6.54 $\pm$ 0.30 7.41
SMaLL-100 4.26 $\pm$ 0.35 7.02 $\pm$ 0.44 4.33 $\pm$ 0.24 9.70 $\pm$ 0.37 4.54 $\pm$ 0.41 5.97
SMaLL-100 (No FT) 5.93 $\pm$ 0.23 6.73 $\pm$ 0.49 6.33 $\pm$ 0.28 10.66 $\pm$ 0.44 5.77 $\pm$ 6.06 7.08
Estimated Model (No FT) 5.76 $\pm$ 0.19 4.29 $\pm$ 0.31 6.23 $\pm$ 0.34 4.59 $\pm$ 0.22 5.67 $\pm$ 0.41 5.31
Ensemble 3.79 $\pm$ 0.30 3.79 $\pm$ 0.25 3.90 $\pm$ 0.22 4.14 $\pm$ 0.24 4.36 $\pm$ 0.22 4.00
Table 1: English-centric test results in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning" and the model inference is done in a zero-shot fashion. We only show M2M100 results for the Unseen setting since NLLB covers all languages in the English-centric dataset. Avg represents the average of the results across the row. The reported results use XGBoost as the regressor. Ensemble denotes combining all four proxy models, the detailed breakdown of this result can be seen in Section B in the Appendix.
Models M2M100 NLLB
LRL MRL LRL MRL
Transformer 6.14 6.67 11.60 10.26
SMaLL-100 4.28 4.55 10.18 7.83
SMaLL-100 (No FT) 6.17 6.96 7.90 11.38
Estimated Model (No FT) 6.15 6.57 4.66 4.32
Ensemble 3.93 3.80 4.10 4.15
Table 2: English-centric LOLO test results in average RMSE by language vitality (lower is better). “No FT" denotes “no fine-tuning". “LRL" and “MRL" denote low-resource languages and medium-resource languages, based on Joshi et al. (2021). The full mapping is in Table 8 & 9 in Section A.1 in the Appendix.
Models Random LOLO
[…]
NLPerf Xia et al. (2020a) with different regressors
XGBoost 2.45 $\pm$ 0.30 1.21 $\pm$ 0.29 7.83 $\pm$ 0.23 8.28 $\pm$ 0.31 4.94
Poly2  Khiu et al. (2024) 4.70 $\pm$ 0.40 4.68 $\pm$ 0.51 7.07 $\pm$ 0.00 7.90 $\pm$ 0.00 6.09
Poly3  Khiu et al. (2024) 4.60 $\pm$ 0.41 4.64 $\pm$ 0.49 7.26 $\pm$ 0.00 8.01 $\pm$ 0.00 6.13
LGBM  Ye et al. (2021a) 2.66 $\pm$ 0.22 1.76 $\pm$ 0.29 7.91 $\pm$ 0.01 8.06 $\pm$ 0.00 5.10
MF Schram et al. (2023) 3.65 $\pm$ 0.26 2.60 $\pm$ 0.39 7.08 $\pm$ 0.23 7.14 $\pm$ 0.22 5.12
ProxyLM (Ours) with different proxy models
Transformer 2.56 $\pm$ 0.43 1.70 $\pm$ 0.20 5.65 $\pm$ 0.23 6.24 $\pm$ 0.34 4.04
SMaLL-100 2.56 $\pm$ 0.33 1.65 $\pm$ 0.44 4.85 $\pm$ 0.36 5.14 $\pm$ 0.46 3.55
SMaLL-100 (No FT) 2.44 $\pm$ 0.21 1.34 $\pm$ 0.38 6.93 $\pm$ 0.34 7.25 $\pm$ 0.37 4.49
Estimated Model (No FT) 2.38 $\pm$ 0.36 1.27 $\pm$ 0.30 5.10 $\pm$ 0.28 5.50 $\pm$ 0.26 3.56
Ensemble 2.41 $\pm$ 0.28 1.56 $\pm$ 0.35 3.73 $\pm$ 0.23 3.79 $\pm$ 0.19 2.90
Table 3: Many-to-Many Languages test results in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning". We do not cover the Unseen setting as most of the languages are not covered by both M2M100 and NLLB in the Many-to-Many Languages dataset. Avg represents the average of the results across the row. The reported results are experiments using XGBoost regressor. Ensemble denotes combining all four proxy models.
3.3 Experimental Settings
Each regressor is evaluated using RMSE as our performance metric and evaluated 5 times. We set our experiment settings as follows:
  • Random: We randomly sample the performance records into training and test sets with a ratio of 7:3. Then, we run 10-fold cross-validation on the training set to find the best hyper-parameters for each regressor. The best-performing regressor would subsequently be evaluated on the test set.
  • Leave-One-Language-Out (LOLO): We select one language as the test set, which is not encountered during training.
  • Unseen: The performance records can be divided into two categories: (1) records with “seen" languages and (2) records with “unseen" languages. “Unseen" languages refer to languages that are not present in the pre-training LM data, while “seen" languages denote those that are present. In this setting, the regressor is trained using records of “seen" languages and tested using records of “unseen" languages.
  • Cross-Dataset: We train the regressor using performance records from the English-centric dataset and test it using records from the Many-to-Many Languages dataset. We opt not to reverse this setup as the dataset exhibits no domain shift and contains fewer performance records.
4 Results and Analysis
In this section, we present the results of the performance predictions for ProxyLM and baselines under the specified settings. Further, we discuss the robustness, effectiveness, and efficiency of ProxyLM in the context of performance prediction.
4.1 English-centric Results
Table 1 shows the overall results on the English-centric dataset. ProxyLM remarkably outperforms all existing baselines. We find that incorporating all proxy models (Ensemble) is the most effective for prediction, leading to a 2.29$\times$ averaged reduction in RMSE across all experimental settings compared to the best baseline. We observe that using the No FT estimated model to predict the performance of their fine-tuned models is surprisingly useful in all settings, especially for NLLB, where the model already has decent machine translation quality on LRLs. This observation is supported by our findings within the XGBoost model that the NLLB No FT feature has the highest importance score among all features, as shown in Figure 16 in the Appendix. Further, using SMaLL-100 fine-tuned performance provides useful estimations for settings involving M2M100 as the estimated model. This may indicate that the performance of a model with similar architecture can be a good estimator for the performance of the larger estimated model. In other words, the choice of proxy model to help prediction matters. Feature importance analysis from the XGBoost model supports this, revealing that the SMaLL-100 fine-tuned feature has the highest importance score among all features, as shown in Figure 15 in the Appendix.
Our analysis also indicates that XGBoost outperforms other regression models, across all evaluated settings. Both XGBoost and LGBM, which are gradient-boosting and tree-based learning methods, demonstrate superior performance metrics across all settings. Their robustness and efficiency as non-linear models are evident when compared to linear models, such as Poly2 and Poly3. Poly2 and Poly3 regressors, which employ second-degree and third-degree polynomial regression approaches respectively, tend to generate lower scores. This diminished performance is largely attributed to their limitations in capturing the nonlinear relationships inherent in the data, leading to suboptimal results. We further present the results by the language vitality in Table 2. The overall difference in RMSE between LRLs and medium-resource languages (MRLs) is relatively small, except when using SMaLL-100 as proxy models. An interesting observation here on SMaLL-100 in predicting the NLLB model is that the No FT model can predict LRLs much better than the fine-tuned counterpart, and the fine-tuned model better predicts MRLs.
[…]
NLPerf Xia et al. (2020a) with different regressors
XGBoost 9.97 $\pm$ 0.35 10.52 $\pm$ 0.73
Poly2  Khiu et al. (2024) 9.27 $\pm$ 0.00 21.43 $\pm$ 0.00
Poly3  Khiu et al. (2024) 8.58 $\pm$ 0.00 19.81 $\pm$ 0.00
LGBM  Ye et al. (2021a) 9.24 $\pm$ 0.24 15.19 $\pm$ 1.00
ProxyLM (Ours) with different proxy models
Transformer 6.65 $\pm$ 0.26 10.69 $\pm$ 0.90
SMaLL-100 5.53 $\pm$ 0.98 7.68 $\pm$ 2.11
SMaLL-100 (No FT) 9.08 $\pm$ 0.63 16.66 $\pm$ 1.08
Estimated Model (No FT) 9.45 $\pm$ 0.78 7.10 $\pm$ 0.69
Ensemble 4.30 $\pm$ 0.17 6.21 $\pm$ 0.50
Table 4: Cross-Dataset test results in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning". The reported results are experiments using the LGBM regressor. Ensemble denotes combining all four proxy models.
4.2 Many-to-Many Languages Results
Table 3 presents the performance of different models on the Many-to-Many Languages dataset. The results reveal that the Ensemble model achieves the lowest RMSE, with a 1.70$\times$ averaged reduction in RMSE across all experimental settings compared to the best baseline, indicating superior accuracy in performance predictions. An exception occurs in the random NLLB setting, where the model utilizing only NLPerf features outperforms the ensemble model, achieving the best performance. Note that no domain shift occurs within the dataset.
A comparative analysis shows that predicting the performance of the M2M100 model in the random setting presents a greater challenge compared to predicting the NLLB model. This discrepancy suggests that the complexity of performance prediction can vary substantially depending on the specific LM and the conditions under which it is evaluated. A particularly noteworthy finding is the effectiveness of using No FT models for estimating LM performance. The No FT models, which do not require any additional fine-tuning, demonstrate high accuracy in their performance predictions. This method offers substantial efficiency benefits, as it eliminates the need for extensive computational resources typically required for model training. In contrast, we find similar results between the LOLO setting for Many-to-Many Languages and English-centric results, where ProxyLM using Ensemble remarkably outperforms all existing baselines. In addition, we find that using SMaLL-100 fine-tuned performance results in better predictions compared to those of the No FT estimated model.
4.3 Cross-Dataset Results
Table 4 illustrates model performance in the Cross-Dataset setup, showcasing the superior performance of ProxyLM with LGBM. We use LGBM instead of XGBoost because LGBM performs better. The results highlight that ProxyLM with Ensemble significantly reduces RMSE compared to the best baseline by 2$\times$ and 1.69$\times$ for M2M100 and NLLB, respectively. This displays consistent performance across datasets and languages that are not encountered during the regressor’s training, including unseen languages for the pre-trained LMs. However, the No FT models exhibit inconsistency compared to other proxy models. We also find varied performance between M2M100 and NLLB, which may be attributed to the English-centric dataset containing only seen languages for NLLB and lacking unseen language examples for the regressor. This highlights the significance of incorporating unseen language instances in training for more robust predictions.
(a) English-centric test results.
(b) Many-to-Many Languages test results.
Figure 2: Ablation study on the LOLO setting with XGBoost on English-centric and Many-to-Many Languages datasets. Proxy Models here indicates Ensemble, which is a combination of all proxy models. Proxy Models significantly reduce RMSE across all scenarios.
Figure 3: RMSE decreases as XGBoost’s training set size increases. Training set size refers to the number of English-centric past performance records.
4.4 Ablation Study
Figure 2 highlights the impact of features used in ProxyLM in the LOLO setting with XGBoost. Utilizing proxy models as features leads to a significant reduction in RMSE across all scenarios, showcasing their importance compared to other features. For the English-centric dataset, including language and dataset features alongside proxy models enhances performance. Dataset features alone show better improvement than language features alone, but the combination of both yields the best performance. On the other hand, for the Many-to-Many Languages dataset, the benefits of incorporating dataset and language features are less pronounced, especially for the M2M100 model, and there may even be a performance dip for the NLLB model due to the dataset’s lack of domain shift.
4.5 Diminishing Returns with Increasing Training Set Size
In Figure 3, we examine the training of the XGBoost regressor using different numbers of English-centric past performance records as the training dataset. While the regressor’s performance shows enhancement with an expanding training size, the incremental benefits start diminishing once the training set surpasses about 400 past performance records. This observation implies that, across datasets, there exists a threshold where the advantages of incorporating additional past performance records begin to exhibit diminishing returns.
Datasets Inference Fine-tuning
English-centric Many-to-Many Langs.
Estimated models
M2M100 338 s 3.94 hrs (4.77$\times$) 1.42 hrs (7.32$\times$)
NLLB 492 s 9.4 hrs (11.48$\times$) 7.21 hrs (37.08$\times$)
Proxy models
SMaLL-100 201 s 2.5 hrs (3.03$\times$) 1.03 hrs (5.29$\times$)
Transformer 60 s 0.83 hrs (1$\times$) 0.19 hrs (1$\times$)
Table 5: Comparison of LMs’ inference time (in seconds) and fine-tuning time (in hours) for one experimental run. The multiplier of fine-tuning time is relative to the Transformer model. All times were calculated using the interquartile mean to ignore outliers. Since the inference time for English-centric and Many-to-Many Languages datasets are roughly the same, only one column is provided.
Regressors English-centric Many-to-Many Langs.
XGBoost 129.29 s 69.81 s
Poly2 0.06 s 0.03 s
Poly3 0.07 s 0.06 s
LGBM 702.12 s 173.17 s
MF N/A 465.37 s
Table 6: Regressor models training time comparison in seconds (s).
Figure 4: Detailed results of XGBoost with ProxyLM Ensemble on M2M100 model under the LOLO setting using the English-centric dataset from Table 1. The results are grouped by (a) Joshi Class and (b) language family that follows the mapping which is provided in Appendix A.1; (c) shows the scatter plot illustrating the correlation of spBLEU scores between the ProxyLM’s prediction and estimated LM, with the light gray dashed line representing the line of equality ($y=x$) with $R^{2}=0.88$ and black dashed line representing Locally Weighted Scatterplot Smoothing (LOWESS) curve to represent the trend.
4.6 Time Efficiency
Table 5 compares the fine-tuning and inference times required for the estimated and proxy models. The results demonstrate that fine-tuning proxy models or direct inference from any model is remarkably faster than fine-tuning all estimated models. Table 6 further illustrates this point, showing only a minimal trade-off in the time needed to train the regressor models. This additional training time is relatively negligible, highlighting the efficiency of using proxy models.
4.7 Performance by Language Categories
In Figure 4, we present detailed XGBoost results with ProxyLM Ensemble on the M2M100 model under the English-centric LOLO experiment, grouped by language categories. Based on the Locally Weighted Scatterplot Smoothing (LOWESS)  Cleveland (1979) curve depicted in Figure 4(c), our method consistently maintains unbiased predictions for spBLEU scores below 40 across various language types. However, as the spBLEU score increases, the availability of data points diminishes, leading to our method under-predicting the performance compared to the true spBLEU score. Outliers observed in Kartvelian languages and Indo-European languages with Joshi class 3 may have contributed to this discrepancy in prediction. These observations suggest that increasing the number of data points covering higher spBLEU scores may help mitigate the bias in prediction. Further experiment details are available in Appendix B.
5 Related Work
The prediction performance of machine learning algorithms has been mainly explored in two research directions: (1) predict the model performance during the training runtime, and (2) predict the model performance by providing extracted features from the dataset Xia et al. (2020a).
Performance Prediction During the Training Runtime
The former aims to infer and extrapolate the learning curve to approximate training results using evaluation metric measurements Kolachina et al. (2012). Domhan et al. (2015) study the quick detection of poor hyper-parameters in probabilistic models after a few steps of Stochastic Gradient Descent (SGD). Adriaensen et al. (2024) extrapolate learning curves from a parametric prior using Markov Chain Monte Carlo (MCMC).
Performance Prediction Using Extracted Features
The latter aims to predict the model performance by learning a correlation between input features and the final evaluation metric. Birch et al. (2008a) identify strong predictive features such as the amount of reordering, the morphological complexity of the target language, and the historical relatedness of the two languages. Xia et al. (2020a) leverage extracted dataset features and typological database language representations. Ye et al. (2021a) introduce the use of confidence intervals and calibration with various regressor algorithms for reliable performance prediction. Schram et al. (2023) apply Bayesian matrix factorization for performance prediction on multilingual NLP tasks. In this work, we focus on exploring the latter. Existing approaches have shown promise using linear regression and gradient-boosting trees Birch et al. (2008b); Xia et al. (2020b); Srinivasan et al. (2021); Ye et al. (2021b). These studies have considered data size, typological features, and language similarity as factors contributing to the model performance.
6 Conclusion
In this paper, we introduce ProxyLM, a novel framework designed to predict the performance of LMs by leveraging proxy models including for LRLs. By utilizing proxy models as substitutes to estimate the performance of the target model, we strategically employ smaller LMs and off-the-shelf models without additional fine-tuning. This framework is highly scalable to multiple proxy models and is task-agnostic, making it applicable to a wide range of downstream NLP tasks. Our streamlined approach showcases substantial advancements in prediction accuracy compared to standard baselines and exhibits strong generalization capabilities across varied scenarios.
Limitations
This paper focuses exclusively on two estimated models: M2M100 and NLLB, to evaluate our proposed framework. For demonstration purposes, we concentrate on the usage of specific models, namely the Transformer model, the SMaLL-100 model, and the No FT models, to illustrate the effectiveness of our proxy models. The M2M100 and NLLB models were selected due to their prominence and relevance in the field of multilingual translation tasks. These models serve as robust benchmarks for assessing the performance and reliability of our proxy-based framework. By using these well-regarded models, we aim to provide compelling evidence of the capabilities and advantages of ProxyLM. While our proposed framework is evaluated solely within the context of machine translation, it is not confined to this application alone. The framework is designed to be versatile and can be extended to a variety of other downstream tasks. We plan to explore these additional applications in future work.
Some other possible avenues for future work could involve a deeper investigation into which proxy models are more effective for enhancing performance prediction in specific settings. Our findings suggest that one proxy model can outperform another in different scenarios, making it crucial to carefully select the most relevant proxy models to maximize the benefits of our approach. Additionally, developing methodologies for collecting relevant past performance records could provide better insights and improve the generalization and accuracy of our framework. Past performance records may provide better information gain than others, potentially minimizing the number of performance records required for a more robust and accurate predictor.
Acknowledgements
We extend our sincere gratitude to Viktoria Schram for providing assistance to reproduce baselines.
Appendix A Experimental Details
Datasets Languages Under Study Domain
English-centric Dataset
   FLoRes-200 Costa-jussà et al. (2022b) afr, amh, arz, bak, bel, ceb, cym, dik, ewe, fao, Multi-domain
guj, hau, hne, hye, ibo, ind, jav, kan, kat, kaz,
khm, kin, kir, lmo, ltz, mar, mri, mya, oci, pan,
plt, sin, sna, snd, som, ssw, tam, tat, tgl, tuk,
wol, xho, yor, zul
   MT560 Gowda et al. (2021)
      Joshua Indian Corpus tam Wiki
      Neulab TED Talk Qi et al. (2018) bel, hye, kat, kaz, mar, mya, tam TED
      News Commentary kaz Economics
      OPUS100 Zhang et al. (2020) afr, amh, bel, cym, guj, hau, hye, ibo, ind, kan, Multi-domain
kat, kaz, khm, kin, kir, mar, mya, oci, pan, sin,
tam, tat, tuk, xho, yor, zul
      OPUS Bible afr, amh, ceb, dik, ewe, guj, hye, ind, kan, mar, Religion
      Christodouloupoulos and Steedman (2015) mri, mya, plt, sin, sna, som, ssw, tam, tgl, wol,
xho, zul
      OPUS OpenSubtitles afr, hye, kat, sin, tgl Movies
      OPUS Tatoeba Tiedemann (2012) afr, amh, arz, bak, bel, ceb, fao, hau, hye, ibo, Conversational
ind, jav, kan, kat, kaz, khm, kin, kir, lmo, ltz,
mar, mri, mya, pan, sna, tat, tgl, tuk, yor, zul
      OPUS Tanzil Tiedemann (2012) amh, hau, ind, snd, som, tat Religion
      OPUS Gnome Tiedemann (2012) fao, mri, som Technology
      OPUS GlobalVoices Tiedemann (2012) amh, mya News
      OPUS EU Bookshop Tiedemann (2012) cym Government
      OPUS SPC afr Government
      OPUS Memat xho Medicine
      OPUS XhosaNavy xho Government
      OPUS KDE4 hne Technology
      OPUS infopankki som Immigration
      PMIndia Haddow and Kirefu (2020) guj, kan, mar, pan, sin, tam Government
      Wiki Titles guj, kaz, tam Wiki
      WikiMatrix Schwenk et al. (2021) arz, bak, bel, ceb, fao, ind, jav, kat, kaz, lmo, Wiki
ltz, mar, oci, sin, tam, tat, tgl
Many-to-Many Languages Dataset
   NusaTranslation Cahyawijaya et al. (2023) bew, btk, ind, jav, mad, mak, min, sun Social Media
Table 7: List of datasets under study covering 50 different languages. We only opt for 44 out of 500 languages available in the MT560 dataset, 44 out of 200 languages available in the FLoRes-200 dataset, and 8 out of 12 languages available in the NusaTranslation dataset.
A.1 Languages Under Study
We list all the languages used in the training from the MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023) datasets in Table 8 and Table 9, respectively. The language code follows ISO639-3 coding. All languages are also complemented by their rarity taxonomy based on Joshi et al. (2021) into two vitality classes: 0-2→low resource language (LRL), and 3→mid resource language (MRL). We also provide information about whether the language was part of the pre-trained M2M100 model dataset to highlight the model knowledge coverage.
A.2 Models
Here are the details on the proxy LMs we use in the experiments as follows:
  • Transformer (100M) Vaswani et al. (2017): a standard encoder-decoder transformer-based model with 6 encoder layers and 6 decoder layers with an embedding dimension of 512. We train the model from randomly initialized parameters with the training set.
  • SMaLL-100 (330M) Mohammadshahi et al. (2022): [^2] SMaLL-100 (330M) is taken from https://github.com/alirezamshi/small100. a distilled version of the M2M100 (12B) model. We utilize the model in two ways: fine-tuned on training data and zero-shot inference.
  • M2M100 (No FT) Fan et al. (2021): [^3] M2M100 (1.2B) is taken from https://github.com/facebookresearch/fairseq/tree/main/examples/m2m_100. a pre-trained estimated model of M2M100 (1.2B) without any fine-tuning. We run the model in a zero-shot fashion.
  • NLLB (No FT) Costa-jussà et al. (2022b): [^4] NLLB (1.3B) is taken from https://github.com/facebookresearch/fairseq/tree/nllb. a pre-trained estimated model of NLLB-200 Distilled (1.3B) without any fine-tuning. We run the model in a zero-shot fashion.
Language Language Code Family Joshi Class Vitality Seen by M2M100
Afrikaans afr indo-european 3 MRL
Amharic amh afro-asiatic 2 LRL
Armenian hye indo-european 1 LRL
Bashkir bak turkic 1 LRL
Belarusian bel indo-european 3 MRL
Burmese mya sino-tibetan 1 LRL
Cebuano ceb austronesian 3 MRL
Chhattisgarhi hne indo-european 0 LRL
Dinka dik nilo-saharan 1 LRL
Egyptian Arabic arz afro-asiatic 3 MRL
Ewe ewe niger-congo 1 LRL
Faroese fao indo-european 1 LRL
Georgian kat kartvelian 3 MRL
Gujarati guj indo-european 1 LRL
Hausa hau afro-asiatic 2 LRL
Igbo ibo niger-congo 1 LRL
Indonesian ind austronesian 3 MRL
Javanese jav austronesian 1 LRL
Kannada kan dravidian 1 LRL
Kazakh kaz turkic 3 MRL
Khmer khm austro-asiatic 1 LRL
Kirghiz kir turkic 1 LRL
Kinyarwanda kin niger-congo 1 LRL
Lombard lmo indo-european 1 LRL
Luxembourgish ltz indo-european 1 LRL
Malagasy plt austronesian 1 LRL
Maori mri austronesian 1 LRL
Marathi mar indo-european 2 LRL
Occitan oci indo-european 1 LRL
Punjabi pan indo-european 2 LRL
Shona sna niger-congo 1 LRL
Sindhi snd indo-european 1 LRL
Sinhala sin indo-european 0 LRL
Somali som afro-asiatic 1 LRL
Swati ssw niger-congo 1 LRL
Tagalog tgl austronesian 3 MRL
Tamil tam dravidian 3 MRL
Tatar tat turkic 1 LRL
Turkmen tuk turkic 1 LRL
Welsh cym indo-european 1 LRL
Wolof wol niger-congo 2 LRL
Xhosa xho niger-congo 2 LRL
Yoruba yor niger-congo 2 LRL
Zulu zul niger-congo 2 LRL
Table 8: List of languages from the English-centric dataset, including their rarity category mapping and an indication of whether they are involved in the pre-training process for M2M100. Note that all languages are involved in the pre-training process for NLLB.
A.3 Hyper-parameters
LM
Each fine-tuning and evaluation for LMs is done with an NVIDIA Tesla V100 32GB GPU. The hyper-parameters used during fine-tuning from the English-centric and Many-to-Many Languages datasets are listed in Table 10, 11, 12, and 13 for SMaLL100, M2M100, NLLB, and Transformer models, respectively.
Regressor
Each regressor is trained on an AMD Ryzen Threadripper 2990WX with 128 GB of RAM and 16 threads. Regressors’ hyper-parameters used are provided in Table 14, 15, 16, and 17 for XGBoost, Poly2/Poly3, LGBM, and MF, respectively. These hyper-parameters were obtained based on the best cross-validation RMSE score using 10 folds.
Language Language Code Family Joshi Class Vitality Seen by M2M100 Seen by NLLB
Indonesian ind austronesian 3 MRL
Javanese jav austronesian 1 LRL
Betawi bew creole 0 LRL
Batak btk austronesian 0 LRL
Madurese mad austronesian 0 LRL
Makassarese mak austronesian 0 LRL
Minangkabau min austronesian 0 LRL
Sundanese sun austronesian 1 LRL
Table 9: List of languages from the Many-to-Many Languages dataset along with their rarity category mapping and an indication of whether they are included in the pre-training process for each respective model.
Hyper-parameter English-centric Many-to-Many Langs.
Encoder Layers 12 12
Decoder Layers 3 3
Encoder Embed Dim 1024 1024
Decoder Embed Dim 1024 1024
Encoder FFN Embed Dim 4096 4096
Decoder FFN Embed Dim 4096 4096
Encoder Attention Heads 16 16
Decoder Attention Heads 16 16
Encoder Layerdrop 0.05 0.05
Decoder Layerdrop 0.05 0.05
Optimizer Adam Adam
Adam Eps 1e-6 1e-6
Adam Betas (0.9, 0.98) (0.9, 0.98)
Patience 6 6
Batch Size 16 16
Dropout 0.1 0.1
Attention Dropout 0.1 0.1
ReLU Dropout 0.1 0.1
Weight Decay 0.0 0.0
Label Smoothing 0.1 0.1
Clip Norm 1.0 1.0
Learning Rate 0.0001 0.0003
Max Tokens (per GPU) 1,000 1,000
Table 10: List of hyper-parameters used for SMaLL100 with English-centric and Many-to-Many Languages datasets.
Hyper-parameter English-centric Many-to-Many Langs.
Encoder Layers 24 24
Decoder Layers 24 24
Encoder Embed Dim 1024 1024
Decoder Embed Dim 1024 1024
Encoder FFN Embed Dim 8192 8192
Decoder FFN Embed Dim 8192 8192
Encoder Attention Heads 16 16
Decoder Attention Heads 16 16
Encoder Layerdrop 0.05 0.05
Decoder Layerdrop 0.05 0.05
Optimizer Adam Adam
Adam Eps 1e-6 1e-6
Adam Betas (0.9, 0.98) (0.9, 0.98)
Patience 6 6
Batch Size 32 32
Dropout 0.1 0.1
Attention Dropout 0.1 0.1
ReLU Dropout 0.1 0.1
Weight Decay 0.0 0.0
Label Smoothing 0.1 0.1
Clip Norm 0.0 0.0
Learning Rate 0.0002 0.0002
Max Tokens (per GPU) 1,792 1,792
Table 11: List of hyper-parameters used for M2M100 with English-centric and Many-to-Many Languages datasets.
A.4 Regressor Dataset Sizes
We provide the details of the regressor’s training and test set size in Table 18.
Appendix B More Detailed Results
We provide detailed visualizations of the results of XGBoost with ProxyLM Ensemble based on multiple language groupings in Figures 5-10 for English-centric result, and Figures 11-14 for Many-to-Many Languages result. Each language groupings plot comprises multiple subplots, including (a) vitality class, (b) Joshi class, (c) language family, and (d) individual languages. The mapping of vitality, Joshi class, and language family follows the classifications in Table 8 and 9.
Appendix C Feature Importances
We provide feature importance scores of XGBoost with ProxyLM Ensemble for the random English-centric experiment in Figure 15 and 16. Each combination consists of one most influential feature followed by others with marginal contributions to the model, each with an importance score of 0.12 or less. We observe that proxy models are always the most influential features in prediction.
Appendix D License for Artifacts
We discuss the license or terms for the use of any artifacts we use in Table 19.
Hyper-parameter English-centric Many-to-Many Langs.
Encoder Layers 24 24
Decoder Layers 24 24
Encoder Embed Dim 1024 1024
Decoder Embed Dim 1024 1024
Encoder FFN Embed Dim 8192 8192
Decoder FFN Embed Dim 8192 8192
Encoder Attention Heads 16 16
Decoder Attention Heads 16 16
Encoder Layerdrop 0.05 0.05
Decoder Layerdrop 0.05 0.05
Optimizer Adam Adam
Adam Eps 1e-6 1e-6
Adam Betas (0.9, 0.98) (0.9, 0.98)
Patience 6 6
Batch Size 32 32
Dropout 0.1 0.1
Attention Dropout 0.1 0.1
ReLU Dropout 0.0 0.0
Weight Decay 0.01 0.01
Label Smoothing 0.1 0.1
Clip Norm 1.0 1.0
Learning Rate 0.00002 0.0001
Max Tokens (per GPU) 1,000 1,000
Table 12: List of hyper-parameters used for NLLB with English-centric and Many-to-Many Languages datasets.
Hyper-parameter English-centric Many-to-Many Langs.
Encoder Layers 6 6
Decoder Layers 6 6
Encoder Embed Dim 512 512
Decoder Embed Dim 512 512
Encoder FFN Embed Dim 2048 2048
Decoder FFN Embed Dim 2048 2048
Encoder Attention Heads 8 8
Decoder Attention Heads 8 8
Encoder Layerdrop 0.05 0.05
Decoder Layerdrop 0.05 0.05
Optimizer Adam Adam
Adam Eps 1e-6 1e-6
Adam Betas (0.9, 0.98) (0.9, 0.98)
Patience 6 6
Batch Size 32 32
Dropout 0.1 0.1
Attention Dropout 0.1 0.1
ReLU Dropout 0.1 0.1
Weight Decay 0.0001 0.0001
Label Smoothing 0.1 0.1
Clip Norm 0 0
Learning Rate 0.001 0.0005
Max Tokens (per GPU) 1,000 1,000
Table 13: List of hyper-parameters used for Transformer with English-centric and Many-to-Many Languages datasets.
Hyper-parameter Many-to-Many Langs. English-centric
M2M100 NLLB M2M100 NLLB
max n_estimators 2000 2000 5000 5000
max eta 0.1 0.1 0.1 0.1
min_child_weight 5 2.5 3 3.5
max_depth 3 3 5 4
gamma 0 0 0.2 0.1
subsample 0.7 0.9 0.95 0.8
colsample_bytree 0.6 0.6 0.85 1.0
reg_alpha 0 0 0.1 0.1
reg_lambda 0.35 0.15 0.1 0.3
Table 14: List of hyper-parameters used for XGBoost Regressor with M2M100 and NLLB models trained with English-centric and Many-to-Many Languages datasets.
Hyper-parameter Value
alpha 0.1
l1_ratio 0.9
Table 15: List of hyper-parameters used for Poly2/Poly3 Regressor with M2M100 and NLLB models trained with English-centric and Many-to-Many Languages datasets.
Hyper-parameter Value
max learning_rate 0.3
max num_leaves 64
n_estimators 100
max_bin 200000
max_depth 10
min_child_weight 0.001
min_child_samples 20
min_split_gain 0.0
colsample_bytree 1.0
subsample 1.0
reg_alpha 0.1
reg_lambda 0.1
Table 16: List of hyper-parameters used for LGBM Regressor with M2M100 and NLLB models trained with English-centric and Many-to-Many Languages datasets. “Max" indicates the maximum value set for the hyper-parameter during the hyper-parameter search.
Hyper-parameter Specification
max alpha 0.01
beta_w 0.1
beta_h 0.1
beta_z 0.01
beta_s 0.01
beta_t 0.01
lr_decay 0.001
iterations 2000
Table 17: List of hyper-parameters used for MF Regressor with M2M100 and NLLB models trained with English-centric and Many-to-Many Languages datasets. “Max" indicates the maximum value set for the hyper-parameter during the hyper-parameter search.
Experimental Settings Train Size Test Size
Random (English-centric) 829 356
Random (Many-to-Many Langs.) 156 68
Unseen 1,084 101
Cross-Dataset 1,185 224
Table 18: Regressor’s training and test set size on different experimental settings. The total experimental records for English-centric and Many-to-Many Languages datasets are 1,185 and 224, respectively.
Datasets URL Link License
MT560 Gowda et al. (2021) https://opus.nlpl.eu/MT560 Unknown
FLoRes Costa-jussà et al. (2022b) Muennighoff/flores200 CC-BY-SA 4.0
NusaTranslation Cahyawijaya et al. (2023) https://huggingface.co/datasets/indonlp/nusatranslation_mt Apache 2.0
Table 19: List of datasets under study with their licenses.
Figure 5: Detailed results of XGBoost with ProxyLM Ensemble on NLLB model under the LOLO setting using the English-centric dataset from Table 1. The results are grouped by (a) Joshi Class and (b) language family that follows the mapping which is provided in Appendix A.1; (c) shows the scatter plot illustrating the correlation of spBLEU scores between the ProxyLM’s prediction and estimated LM, with the gray dashed line representing the line of equality ($y=x$) with $R^{2}=0.89$ and black dashed line representing Locally Weighted Scatterplot Smoothing (LOWESS) curve to represent the trend.
Figure 6: Detailed results of XGBoost with ProxyLM Ensemble under the LOLO setting using the English-centric dataset from Table 1, grouped by vitality, for (a) M2M100 and (b) NLLB model. Vitality is outlined in Table 8.
Figure 7: Detailed results of XGBoost with ProxyLM Ensemble on the M2M100 model under the LOLO setting using the English-centric dataset from Table 1, categorized by language.
Figure 8: Performance heatmap of XGBoost with ProxyLM Ensemble on the M2M100 model under the LOLO setting using the English-centric dataset from Table 1. The heatmap showcases the performance based on vitality level and language family, as outlined in Table 8.
Figure 9: Detailed results of XGBoost with ProxyLM Ensemble on the NLLB model under the LOLO setting using the English-centric dataset from Table 1, categorized by language.
Figure 10: Performance heatmap of XGBoost with ProxyLM Ensemble on the NLLB model under the LOLO setting using the English-centric dataset from Table 1. The heatmap illustrates the model’s performance based on vitality levels and language family classifications as specified in Table 8.
Figure 11: Detailed results of XGBoost with ProxyLM Ensemble on the M2M100 model under the LOLO setting using the Many-to-Many Languages dataset from Table 3, organized by language groupings. The mapping of (a) Vitality, (b) Joshi Class, and (c) language family align with the classifications provided in Table 9.
Figure 12: Detailed results of XGBoost with ProxyLM Ensemble on the M2M100 model under the LOLO setting using the Many-to-Many Languages dataset from Table 3, categorized by language.
Figure 13: Detailed results of XGBoost with ProxyLM Ensemble on the NLLB model under the LOLO setting using the Many-to-Many Languages dataset from Table 3, organized by language groups. The mapping of (a) Vitality, (b) Joshi Class, and (c) language family corresponds to the classifications provided in Table 9.
Figure 14: Detailed results of XGBoost with ProxyLM Ensemble on the NLLB model under the LOLO setting using the Many-to-Many Languages dataset from Table 3, categorized by language.
Figure 15: Feature importance analysis of XGBoost with ProxyLM Ensemble on the M2M100 model using the English-centric dataset.
Figure 16: Feature importance analysis of XGBoost with ProxyLM Ensemble on the NLLB model using the English-centric dataset.