### v3

# ProxyLM: Predicting Language Model Performance on Multilingual Tasks via Proxy Models

Abstract
Performance prediction is a method to estimate the performance of Language Models (LMs) on various Natural Language Processing (NLP) tasks, mitigating computational costs associated with model capacity and data for fine-tuning. Our paper presents ProxyLM, a scalable task- and language-agnostic framework designed to predict the performance of LMs using proxy models. These proxy models act as surrogates, approximating the performance of the LM of interest. By leveraging these proxy models, ProxyLM significantly reduces computational overhead in task evaluations, achieving up to a 37.08$\times$ speedup over traditional methods, even with our smallest proxy models. Our results across multiple multilingual NLP tasks and various robustness tests demonstrate that ProxyLM not only adapts well to previously unseen languages in pre-trained LMs, but also generalizes effectively across different datasets, outperforming the state-of-the-art by at least 1.78$\times$ in terms of root-mean-square error (RMSE).
ProxyLM: Predicting Language Model Performance on Multilingual
Tasks via Proxy Models
David Anugraha1, Genta Indra Winata2thanks: The work was conducted outside Capital One., Chenyue Li1, Patrick Amadeus Irawan3, En-Shiun Annie Lee1,4 1University of Toronto  2Capital One 3Institut Teknologi Bandung  4Ontario Tech University anugraha@cs.toronto.edu, genta.winata@capitalone.com, {jimschenchen,patrickamadeusirawan}@gmail.com, annie.lee@ontariotechu.ca
1 Introduction
Language Models (LMs) have become increasingly valuable for assessing Natural Language Processing (NLP) tasks Raffel et al. (2020); Brown et al. (2020); Costa-jussà et al. (2022); Touvron et al. (2023a, b); Workshop et al. (2022). However, fine-tuning and evaluating these models are resource-intensive processes in terms of both computation and time. These costs escalate with model size, especially when experimenting across multiple datasets. As highlighted in Kaplan et al. (2020), there is a scaling law that applies to both model and dataset sizes, and computational demands, indicating that larger models and broader datasets require increased computational resources. Modeling low-resource languages (LRLs) in multilingual contexts presents a range of challenges. One significant challenge is the limited data availability, which hampers effective fine-tuning processes Gu et al. (2018); Adilazuarda et al. (2024), making model adaptation through fine-tuning a challenging task Zoph et al. (2016); Liu et al. (2021). Another critical issue is the lack of pre-training data for numerous regional languages, such as Southeast Asian languages Winata et al. (2022, 2024b); Yong et al. (2024), with many languages being omitted during the pre-training phase of multilingual LMs.
Figure 1: ProxyLM framework for LM performance prediction. (Top) The evaluation metric is computed on the test set using a proxy model $\mathcal{M}_{p}^{i}$. (Bottom) The regressor $g$ is trained using proxy model scores as well as dataset and language features by minimizing the RMSE difference of $y_{\mathcal{M}}$ and $\hat{y}_{\mathcal{M}}$.
Given the limited academic computational resources for LM fine-tuning and inadequate datasets for LRLs, performance prediction is an efficient method to alleviate the dependency on extensive resources by leveraging past performance records on an NLP task. While linear regression and gradient-boosting hold promise in performance prediction Birch et al. (2008); Srinivasan et al. (2021); Xia et al. (2020); Ye et al. (2021); Schram et al. (2023); Khiu et al. (2024), existing solutions primarily focus on homogeneous data settings and prioritize high-resource languages using Transformer models Vaswani et al. (2017). Khiu et al. (2024) examine diverse datasets and LRLs but encounter limitations in the number of experiments, language diversity, and model scope, focusing solely on mBART Liu et al. (2020). Recent advancements in larger multilingual models, like NLLB Costa-jussà et al. (2022) and M2M100 Fan et al. (2021), have significantly improved machine translation capabilities, exceeding those of mBART and other LMs Zhu et al. (2023).
In this paper, we propose ProxyLM, [^1] We release our code at https://github.com/davidanugraha/proxylm. a scalable task- and language- agnostic framework to predict LM performance by utilizing proxy models. Proxy models are defined as substitute models, wherein the performance of these substitute models are used to estimate the performance of another LM. This other model can be significantly larger than our proxy models. For optimizing the prediction, we utilize much smaller LMs as proxy models and off-the-shelf models without further tuning. This approach is very scalable to multiple proxy models and task-agnostic to any modalities, thus it can be applied to any downstream tasks. This study focuses on three multilingual tasks, covering machine translation (MT), intent classification, and slot filling. Our approach outperforms the existing work from Xia et al. (2020); Ye et al. (2021); Schram et al. (2023); Khiu et al. (2024), which opens a new avenue to employ LMs for model performance prediction. Therefore the contribution of our paper can be summarized in three-fold:
  1. 1.
    We introduce ProxyLM, an efficient and scalable task- and language-agnostic framework designed to predict the performance of LMs. This framework significantly reduces the computational costs associated with fine-tuning and inference during model selection.
  2. 2.
    We demonstrate the effectiveness and robustness of ProxyLM across 34 dataset sources and 56 languages for MT, and 51 languages for intent classification and slot filling, each on two estimated LMs. Our framework substantially outperforms all existing baselines in all NLP tasks, datasets, and settings, including scenarios involving extremely LRLs that remain unseen by pre-trained LMs and across different datasets, surpassing the state-of-the-art performance measured with root-mean-square error (RMSE) by at least 1.78$\times$.
  3. 3.
    We also provide a time analysis comparing the deployment of proxy models for performance prediction to direct LM fine-tuning. Our results indicate that, with our smallest proxy models, we can achieve up to a 37.08$\times$ speedup on task evaluation compared to the traditional approach, highlighting the efficiency of our approach.
2 Methodology
In this section, we formally define the LM performance prediction problem and our proposal to improve performance prediction.
2.1 ProxyLM
Recall that performance prediction is a task of estimating a system’s performance based on the model and its training strategy, training and test dataset, and language used. Formally, let LM $\mathcal{M}$ be our estimated model. $\mathcal{M}$ is trained over a training dataset $\mathcal{D}$ with source language $\mathcal{L}_{s}$ and target language $\mathcal{L}_{t}$, and then tested using dataset $\mathcal{D}^{\prime}$. $\mathcal{M}$’s performance, denoted $y_{\mathcal{M}}$, can be formulated under function $f$ that relates between these variables:
$y_{\mathcal{M}}=f(\mathcal{M},\mathcal{D},\mathcal{D}^{\prime},\mathcal{L}_{s},\mathcal{L}_{t}).$ (1)
We can approximate $f$ by transforming Equation 1 into a regression task with a regressor function $g$, which will be trained on past performance records. Previous works  Xia et al. (2020); Ye et al. (2021); Schram et al. (2023); Khiu et al. (2024) formulate regressor that takes dataset features $\Phi(\mathcal{D},\mathcal{D}^{\prime})$ to identify the characteristics of the training and test datasets, as well as the distribution shift between them. It also takes language features $\Psi(\mathcal{L}_{s},\mathcal{L}_{t})$ to measure the similarities between the source and target languages. This can be formulated as follows:
$\hat{y_{\mathcal{M}}}=g(\Phi(\mathcal{D},\mathcal{D}^{\prime});\Psi(\mathcal{L}_{s},\mathcal{L}_{t})).$ (2)
We present ProxyLM, a framework that leverages the past performance of other models, referred to as proxy models, as additional context for our regressor. Intuitively, proxy models can provide valuable insights that assist in predicting the performance of the estimated model $\mathcal{M}$, which addresses the gap in previous works for not accounting $\mathcal{M}$. Formally, let $\mathcal{M}_{p}=[\mathcal{M}_{p}^{1},\dots,\mathcal{M}_{p}^{N}]$ be a set of $N$ proxy models. To integrate the information from these proxy models, we modify Equation 2 as follows:
$\hat{y_{\mathcal{M}}}=g(\hat{y}_{\mathcal{M}_{p}};\Phi(\mathcal{D},\mathcal{D}^{\prime});\Psi(\mathcal{L}_{s},\mathcal{L}_{t})),$ (3)
where $y_{\mathcal{M}_{p}}=[y_{\mathcal{M}_{p}^{1}},\dots,y_{\mathcal{M}_{p}^{N}}]$ represents the performance records of $N$ proxy models. The advantage of using proxy models arises from their faster fine-tuning and evaluation compared to the estimated model $\mathcal{M}$. This also means that off-the-shelf models can be used directly without additional tuning if they already perform the task adequately, further enhancing efficiency.
2.2 ProxyLM Features
Language Features.
We use URIEL Typological Database Littell et al. (2017) similar to Xia et al. (2020) including geographic, genetic, inventory, syntactic, phonological, and featural distance. The language features are useful to provide a language-specific representation to the regressor.
Dataset Features.
We extract six features from the dataset, including train size, vocab size, average sentence length, word overlap, Type-Token Ratio (TTR), and TTR distance from $\mathcal{D}$ and $\mathcal{D^{\prime}}$ based on  Xia et al. (2020). We will refer to these features and language features combined as NLPerf features. Furthermore, we incorporate the distribution shift information between the training and test datasets using Jensen-Shannon Divergence (JSD) as described by Khiu et al. (2024). In addition, we include the cosine similarity of term frequency-inverse document frequency (TF-IDF) representations and sentence embeddings using Sentence-BERT Reimers and Gurevych (2019). Details on how these features are computed can be found in Appendix Section A.
Proxy Models Features.
We leverage the performance data from proxy models, derived by averaging results from multiple fine-tuning and evaluation iterations on identical datasets and languages. Moreover, we retain the flexibility to adjust the number of proxy models employed, facilitating efficient and adaptable performance estimation.
3 Experimental Setup
In this section, we describe the datasets and LMs used to obtain LMs’ performance records. These records are then used to train various regressor models under different experimental settings to investigate our approach to performance predictions. The details of the hyper-parameters for both the LMs and the regressors are provided in B.3.
3.1 Datasets
Machine Translation.
We use two types of datasets: English-centric and Many-to-Many Languages. The English-centric dataset involves English serving as either the source or target language. Our English-centric dataset comes from the MT560 Gowda et al. (2021) dataset, where we curate 32 datasets and select 50 languages out of 500 for evaluation. Furthermore, we use the FLoRes-200 dataset Costa-jussà et al. (2022) for additional validation and test sets. These datasets consist of translations with varying quality across diverse domains, presenting significant challenges for a robust performance prediction. In contrast, the Many-to-Many Languages dataset allows any language to act as the source or target. We use the NusaTranslation dataset Cahyawijaya et al. (2023) as our Many-to-Many Languages dataset, which comprises parallel texts in 12 Indonesian regional languages. As many of these languages are absent in pre-trained multilingual models, we analyze 8 out of the 12 languages due to limited data in the remaining 4. Both datasets encompass 56 languages across various domains such as economics, technology, and medicine. Detailed language insights are available in the Appendix Section B.1.
Intent Classification and Slot Filling.
We use MASSIVE FitzGerald et al. (2022) as our dataset encompassing 51 languages. We fine-tune each LM on all languages and evaluate it on all languages. Detailed language insights are available in the Appendix Section B.1.
3.2 Estimated LMs
Machine Translation.
We employ two estimated LMs: M2M100 1.2B Fan et al. (2021) and NLLB 1.3B Costa-jussà et al. (2022). Each estimated model is fine-tuned using a standard next-token prediction objective on the training set.
Intent Classification and Slot Filling.
We employ two decoder-only estimated LMs: LLaMA-3 Instruct (8B)  Dubey et al. (2024) and Aya-23 (8B) Aryabumi et al. (2024). Both models are fine-tuned using supervised fine-tuning (SFT) combined with Low-Rank Adaptation (LoRA) Hu et al. (2021) on the training set.
3.3 Proxy Models
Machine Translation.
We utilize four different transformer-based models: an encoder-decoder random initialized Transformers (100M) Vaswani et al. (2017), SMaLL-100 (330M) Mohammadshahi et al. (2022), M2M100 Fan et al. (2021), and NLLB  Costa-jussà et al. (2022). For M2M100 and NLLB, we use the models without any additional tuning (No FT) in a zero-shot fashion. For simplicity, the term “fine-tuning" will be used throughout this paper to refer to both the process of training from scratch (as in the case of the Transformer (100M) model) and the process of fine-tuning pre-trained LMs. Model details are provided in the Appendix Section  B.2. The evaluation is primarily conducted using SentencePiece BLEU (spBLEU) Goyal et al. (2022), which has proven to be a reliable metric in multilingual and LRLs. For a more comprehensive assessment, we also use COMET-22 Rei et al. (2022), as it shows a high correlation with human judgments Freitag et al. (2023). However, COMET-22 is applied to only a subset of the English-centric dataset due to its limited language coverage, especially to LRLs.
Intent Classification and Slot Filling.
We utilize three decoder-only models: SmolLM (135M and 360M) Allal et al. (2024) and BLOOMZ (560M) Muennighoff et al. (2022). Model details are provided in the Appendix Section  B.2. The evaluation is done using accuracy for intent classification and micro-F1 for slot filling.
3.4 Regressor Models
We utilize XGBoost Chen and Guestrin (2016), LGBM Ye et al. (2021), Poly2 Khiu et al. (2024), and Poly3 Khiu et al. (2024) as our regressors. In most of our experiments, we apply XGBoost as our default regressor because we find it to be the best-performing model based on the cross-validation during training, while the other regressors serve as baselines. Specifically for MT in the Many-to-Many Languages setting, Matrix Factorization with context features (MF) is used as an additional baseline Schram et al. (2023). We do not apply MF to our English-centric setting because MF requires the performance records to be structured in two dimensions—one for the source language and one for the target language. In the English-centric setting, this would result in a sparse matrix with only one fully populated row or column, corresponding to English, making MF impractical for this setup.
English-centric Many-to-Many
Models Random LOLO Avg. Random LOLO Avg.
[…]
NLPerf Xia et al. (2020) with different regressors
XGBoost 7.69 7.73 9.20 12.92 9.39 2.45 1.11 7.83 8.28 4.94
Poly2 Khiu et al. (2024) 11.21 16.23 15.55 43.02 21.50 4.70 4.68 7.07 7.90 6.09
Poly3 Khiu et al. (2024) 11.00 15.64 62.29 236.29 81.31 4.60 4.64 7.26 8.01 6.13
LGBM Ye et al. (2021) 7.88 8.15 9.71 12.81 9.64 3.65 2.60 7.08 7.14 5.12
ProxyLM (Ours) with different proxy models
Transformer 4.68 7.22 6.18 11.78 7.47 2.56 1.70 5.65 6.24 4.04
SMaLL-100 4.07 6.33 4.59 10.33 6.33 2.56 1.65 4.85 5.14 3.55
SMaLL-100 (No FT) 5.27 6.04 6.28 10.94 7.13 2.44 1.34 6.93 7.25 4.49
Estimated Model (No FT) 5.23 4.15 6.18 5.42 5.25 2.38 1.27 5.10 5.50 3.56
Ensemble 3.21 3.68 3.74 4.94 3.89 2.41 1.56 3.73 3.79 2.90
Table 1: MT test results on English-centric and Many-to-Many Languages datasets using spBLEU in average RMSE (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning" and the model inference is done in a zero-shot fashion. Avg represents the average of the results across the row for each respective dataset. The reported results use XGBoost as the regressor. Ensemble denotes combining all four proxy models, the detailed breakdown of this result with the standard deviation can be seen in the Appendix Section C.
3.5 Experimental Settings
Each regressor is evaluated using RMSE as our performance metric and evaluated 5 times. For all tasks, we set our experiment settings as follows:
  • Random: We randomly sample the performance records into training and test sets with a ratio of 7:3. Then, we run 10-fold cross-validation on the training set to find the best hyper-parameters for each regressor. The best-performing regressor would subsequently be evaluated on the test set.
  • Leave-One-Language-Out (LOLO): We select one language as the test set, which is not encountered during training.
To test the robustness of ProxyLM, we also provide two additional setups specifically for MT:
  • Unseen: The performance records for MT can be divided into two categories: (1) records with “seen" languages and (2) records with “unseen" languages. “Unseen" languages refer to languages that are not present in the pre-training LM data, while “seen" languages denote those that are present. In this setting, the regressor is trained using records of “seen" languages and tested using records of “unseen" languages.
  • Cross-Dataset: The regressor can be trained using performance records from the English-centric dataset and tested using the Many-to-Many Languages dataset. We opt not to reverse this setup as the Many-to-Many dataset exhibits no domain shift and contains fewer performance records.
4 Results and Analysis
In this section, we present the results of the performance predictions for ProxyLM and baselines under the specified settings. Further, we discuss the robustness, effectiveness, and efficiency of ProxyLM in the context of performance prediction.
4.1 Machine Translation
4.1.1 English-centric Results
Table 1 shows the overall results on the English-centric dataset using spBLEU. ProxyLM remarkably outperforms all existing baselines. We find that incorporating all proxy models (Ensemble) is the most effective for prediction, leading to a 2.41$\times$ averaged reduction in RMSE across all experimental settings compared to the best baseline. Note that the significant improvement remains consistent when evaluated using a different metric, such as COMET-22, which yields a 2.00× averaged reduction in RMSE across experimental settings compared to the best baseline, as shown in Table 14 in Appendix Section A. We observe that using the No FT estimated model to predict the performance of their fine-tuned models is surprisingly useful in all settings, especially for NLLB, where the model already has decent machine translation quality on LRLs. This observation is supported by our findings within the XGBoost model that the NLLB No FT feature has the highest importance score among all features, as shown in Figure 9 in the Appendix. Furthermore, using SMaLL-100 fine-tuned performance provides useful estimations for settings involving M2M100 as the estimated model. This may indicate that the performance of a model with similar architecture can be a good estimator for the performance of the larger estimated model. In other words, the choice of proxy model to help prediction matters. Feature importance analysis from the XGBoost model supports this, revealing that the SMaLL-100 fine-tuned feature has the highest importance score among all features, as shown in Figure 8 in the Appendix.
Intent Classification Slot Filling
Models Random LOLO Avg. Random LOLO Avg.
Aya23$\downarrow$ LLaMA3$\downarrow$ Aya23$\downarrow$ LLaMA3$\downarrow$ Aya23$\downarrow$ LLaMA3$\downarrow$ Aya23$\downarrow$ LLaMA3$\downarrow$
NLPerf Xia et al. (2020) with different regressors
XGBoost 0.0761 0.0191 0.1573 0.0581 0.0777 0.0693 0.0548 0.1219 0.1093 0.0888
Poly2 Khiu et al. (2024) 0.1996 0.0979 0.2075 0.0918 0.1492 0.1396 0.1412 0.1418 0.1414 0.1410
Poly3 Khiu et al. (2024) 0.1990 0.0969 0.2191 0.0925 0.1519 0.1393 0.1401 0.1448 0.1413 0.1414
LGBM Ye et al. (2021) 0.0839 0.0198 0.1545 0.0558 0.0785 0.0692 0.0557 0.1218 0.1152 0.0905
ProxyLM (Ours) with different proxy models
SmolLM (135M) 0.0676 0.0171 0.1273 0.0455 0.0644 0.0618 0.0538 0.1004 0.0953 0.0778
SmolLM (360M) 0.0604 0.0157 0.1118 0.0441 0.0580 0.0562 0.0506 0.0844 0.0868 0.0695
BLOOMZ (560M) 0.0692 0.0179 0.1283 0.0482 0.0659 0.0618 0.0540 0.1023 0.0995 0.0794
Ensemble 0.0609 0.0164 0.1112 0.0442 0.0582 0.0561 0.0508 0.0830 0.0884 0.0696
Table 2: Intent classification and slot filling results using accuracy and micro-F1 score, respectively, with average RMSE (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. Avg represents the average of the results across the row for each respective task. The reported results use XGBoost as the regressor for both intent classification and slot filling. Ensemble denotes combining all three proxy models. The detailed breakdown of this result with the standard deviation can be seen in the Appendix Section C.
4.1.2 Many-to-Many Languages Results
Table 1 presents the performance of different models on the Many-to-Many Languages dataset. The results reveal that the Ensemble model achieves the lowest RMSE, with a 1.70$\times$ averaged reduction in RMSE across all experimental settings compared to the best baseline, indicating superior accuracy in performance predictions. An exception occurs in the random NLLB setting, where the model utilizing only NLPerf features outperforms the ensemble model, achieving the best performance. Note that no domain shift occurs within the dataset.
A comparative analysis shows that predicting the performance of the M2M100 model in the random setting presents a greater challenge compared to predicting the NLLB model. This discrepancy suggests that the complexity of performance prediction can vary substantially depending on the specific LM and the conditions under which it is evaluated. A particularly noteworthy finding is the effectiveness of using No FT models for estimating LM performance. The No FT models, which do not require any additional fine-tuning, demonstrate high accuracy in their performance predictions. This method offers substantial efficiency benefits, as it eliminates the need for extensive computational resources typically required for model training. In contrast, we find similar results between the LOLO setting for Many-to-Many Languages and English-centric results, where ProxyLM using Ensemble remarkably outperforms all existing baselines. In addition, we find that using SMaLL-100 fine-tuned performance results in better predictions compared to those of the No FT estimated model.
4.2 Intent Classification and Slot Filling
Table 2 presents the overall results for both intent classification and slot filling tasks. In these experiments, ProxyLM employs XGBoost as the regressor for intent classification and LGBM for slot filling, as LGBM exhibited superior performance during training cross-validation for the slot filling task. Consistent with our findings in MT, ProxyLM outperforms all existing baselines in both tasks. Notably, the SmolLM (360M) model emerges as the most effective proxy model, achieving an average RMSE reduction of 1.34$\times$ in intent classification and 1.28$\times$ in slot filling across all experimental settings when compared to the best baseline. As with MT, the choice of proxy model significantly impacts prediction performance. Based on the feature importance analysis in Figure 10-13 in the Appendix, SmolLM (360M) exhibits the highest importance score among all proxy models in both tasks. This suggests that the size of the proxy models may not be a reliable indicator of their effectiveness.
5 Ablation Study
5.1 Robustness of ProxyLM
Figure 2 illustrates the performance of our models in both the Unseen and Cross-Dataset setups, highlighting the robustness results achieved by ProxyLM. For the Cross-Dataset evaluation, we opted for LGBM instead of XGBoost, as it demonstrated better performance during cross-validation on the training set. ProxyLM with Ensemble shows a significant reduction in RMSE compared to the best baseline: a 1.84$\times$ reduction in the Unseen setup, and reductions of 2.15$\times$ and 1.78$\times$ for M2M100 and NLLB in the Cross-Dataset scenario, respectively. This consistent performance across datasets and languages—including those not encountered during the regressor’s training, such as unseen languages for the pre-trained LMs—emphasizes the model’s generalization capabilities. We also observe better performance for M2M100 compared to NLLB, which may be attributed to NLLB’s reliance on an English-centric dataset containing only seen languages, lacking examples of unseen languages for the regressor. This may indicate the importance of including instances of unseen languages in the regressor training dataset for achieving more robust predictions.
Figure 3 shows the impact of features used in ProxyLM for MT in the LOLO setting with XGBoost. Utilizing proxy models as features leads to a significant reduction in RMSE across all scenarios, showcasing their importance compared to other features. For the English-centric dataset, including language and dataset features alongside proxy models enhances performance. Dataset features alone show better improvement than language features alone, but the combination of both yields the best performance. On the other hand, for the Many-to-Many Languages dataset, the benefits of incorporating dataset and language features are less pronounced, especially for the M2M100, and there may even be a performance dip for the NLLB due to the dataset’s lack of domain shift.
Figure 2: Unseen and Cross-Dataset MT test results on English-centric dataset in average RMSE (lower is better). We only show the best-performing baseline for comparison with ProxyLM with different proxy models. “No FT" denotes “no fine-tuning". We only show M2M100 results for the Unseen setting since NLLB covers all languages in the English-centric dataset. The reported results for the Unseen setting use XGBoost, while the Cross-Dataset experiments use LGBM. Ensemble denotes combining all four proxy models. The detailed breakdown of this result with the standard deviation can be seen in the Appendix Section C.
(a) English-centric test results.
(b) Many-to-Many Languages test results.
Figure 3: Ablation study on the LOLO setting with XGBoost on English-centric and Many-to-Many Languages datasets. Proxy Models here indicates Ensemble, which is a combination of all proxy models. Proxy Models significantly reduce RMSE across all scenarios.
Datasets Inference Fine-tuning
English-centric Many-to-Many Langs.
Estimated models
M2M100 421 s 3.94 hrs (7.04$\times$) 1.42 hrs (7.32$\times$)
NLLB 737 s 12.08 hrs (21.57$\times$) 7.21 hrs (37.08$\times$)
Proxy models
SMaLL-100 333 s 2.38 hrs (4.25$\times$) 1.03 hrs (5.29$\times$)
Transformer 231 s 0.56 hrs (1$\times$) 0.19 hrs (1$\times$)
Table 3: Comparison of LMs’ inference time (in seconds) and fine-tuning time (in hours) for one MT experimental run. The multiplier of fine-tuning time is relative to the Transformer model. All times were calculated using the interquartile mean to ignore outliers.
Datasets Intent Classification Slot Filling
Estimated models
Aya-23 0.68 hrs (6.13$\times$) 5.10 hrs (13.48$\times$)
LLaMA3 0.54 hrs (4.91$\times$) 4.36 hrs (11.53$\times$)
Proxy models
BLOOMZ-560M 0.12 hrs (1.06$\times$) 0.59 hrs (1.57$\times$)
SmolLM-360M 0.12 hrs (1.09$\times$) 0.40 hrs (1.05$\times$)
SmolLM-150M 0.11 hrs (1.00$\times$) 0.38 hrs (1.00$\times$)
Table 4: Comparison LMs’ overall fine-tuning and evaluation time (in hours) for one intent classification or slot filling experimental run. The multiplier of the time is relative to SmolLM-135M model as it is the smallest proxy model. All times were calculated using the interquartile mean to ignore outliers.
Figure 4: Detailed results of XGBoost with ProxyLM Ensemble on M2M100 model under the LOLO setting on MT using the English-centric dataset from Table 1. The results are grouped by (a) Joshi Class and (b) language family that follows the mapping which is provided in Appendix B.1; (c) shows the scatter plot illustrating the correlation of spBLEU scores between the ProxyLM’s prediction and estimated LM, with the light gray dashed line representing the line of equality ($y=x$) with $R^{2}=0.90$ and black dashed line representing Locally Weighted Scatterplot Smoothing (LOWESS) curve to represent the trend.
5.2 Time Efficiency
Table 3 compares the fine-tuning and inference times required for the estimated and proxy models on MT, while Table  4 compares the fine-tuning and evaluation times required for the estimated and proxy models on intent classification and slot filling. The results demonstrate that fine-tuning proxy models or direct inference from any model is remarkably faster than fine-tuning all estimated models. Table 5 further illustrates this point, showing only a minimal trade-off in the time needed to train the regressor models. This additional training time is relatively negligible, highlighting the efficiency of using proxy models.
Regressors English-centric (MT) Many-to-Many (MT) Intent and Slot
XGBoost 2.24 s 0.87 s 6.33 s
Poly2 0.07 s 0.06 s 0.09 s
Poly3 0.06 s 0.06 s 0.10 s
LGBM 90.79 s 27.51 s 118.24 s
MF N/A 141.69 s N/A
Table 5: Regressor models training time (in seconds) per one round of cross-validation with 10-folds across all setups. We combine the timing for intent classification and slot filling since they both contain the same amount of training data. All times were calculated using the interquartile mean to ignore outliers.
5.3 Performance by Language Categories
In Figure 4, we present detailed XGBoost results with ProxyLM Ensemble on the M2M100 model under the English-centric LOLO experiment, grouped by language categories. ProxyLM demonstrates relatively stable performance across languages belonging to different Joshi classes and linguistic families. Based on the Locally Weighted Scatterplot Smoothing (LOWESS)  Cleveland (1979) curve depicted in Figure 4(c), our method consistently maintains unbiased predictions for spBLEU scores below 40 across various language types. However, as the spBLEU score increases, the availability of data points diminishes, leading to our method under-predicting the performance compared to the true spBLEU score. Outliers observed in Kartvelian languages and Indo-European languages with Joshi class 3 may have contributed to this discrepancy in prediction. These observations suggest that increasing the number of data points covering higher spBLEU scores may help mitigate the bias in prediction.
6 Related Work
The prediction performance of machine learning algorithms has been mainly explored in two research directions: (1) predict the model performance during the training runtime, and (2) predict the model performance by providing extracted features from the dataset Xia et al. (2020).
Performance Prediction During the Training Runtime.
The former aims to infer and extrapolate the learning curve to approximate training results using evaluation metric measurements Kolachina et al. (2012). Domhan et al. (2015) study the quick detection of poor hyper-parameters in probabilistic models after a few steps of Stochastic Gradient Descent (SGD). Adriaensen et al. (2024) extrapolate learning curves from a parametric prior using Markov Chain Monte Carlo (MCMC).
Performance Prediction Using Extracted Features.
The latter aims to predict the model performance by learning a correlation between input features and the final evaluation metric. Birch et al. (2008) identify strong predictive features such as the amount of reordering, the morphological complexity of the target language, and the historical relatedness of the two languages. Xia et al. (2020) leverage extracted dataset features and typological database language representations. Ye et al. (2021) introduce the use of confidence intervals and calibration with various regressor algorithms for reliable performance prediction. Schram et al. (2023) apply Bayesian matrix factorization for performance prediction on multilingual NLP tasks. In this work, we focus on exploring the latter. Existing approaches have shown promise using linear regression and gradient-boosting trees Birch et al. (2008); Xia et al. (2020); Srinivasan et al. (2021); Ye et al. (2021). These studies have considered data size, typological features, and language similarity as factors contributing to the model performance.
Enhancing LLMs through Small Models.
Recent studies, as compiled in Chen and Varoquaux (2024), have explored leveraging smaller models to complement LLMs across various tasks. A closely related application involves using smaller models to assist with LLM inference and evaluation Kuhn et al. (2023); Manakul et al. (2023); Wang et al. (2024); Winata et al. (2024a). In contrast, our work focuses on employing proxy models within a language- and task-agnostic framework to predict LLM performance accurately. This approach offers a cost-effective alternative to fine-tuning and inference during model selection, while achieving high evaluation accuracy and robustness across extreme domain shifts. These include datasets spanning diverse domains, quality levels, and languages, with particular effectiveness demonstrated for extremely LRLs.
7 Conclusion
In this paper, we introduce ProxyLM, a novel framework designed to predict the performance of LMs by leveraging proxy models including for LRLs. By utilizing proxy models as substitutes to estimate the performance of the target model, we strategically employ smaller LMs or off-the-shelf models without additional fine-tuning. This framework is highly scalable to multiple proxy models and is task- and language-agnostic, making it applicable to a wide range of downstream NLP tasks. Our approach showcases substantial advancements in prediction accuracy compared to standard baselines and exhibits strong generalization capabilities across varied scenarios.
Limitations
This paper focuses on the empirical use of various proxy models without delving into the intricacies of the proxy model selection. Specifically, we do not investigate methods for determining which proxy models are the most effective across different contexts without relying on empirical experimentation. This limitation highlights a significant avenue for future research, which could involve a more systematic approach to identifying the most effective proxy models given estimated LMs and NLP tasks. Nevertheless, the use of proxy models consistently outperforms the absence of such models, demonstrating their importance in performance prediction.
Alternatively, developing robust methodologies for collecting and analyzing relevant past performance records could provide invaluable insights that enhance the generalization and accuracy of our predictive framework. Some performance data may offer greater information gain than others, potentially minimizing the number of performance records required to achieve a more robust and accurate predictor. By establishing a clearer understanding of which performance records yield the most informative insights, we can optimize our approach and improve our overall predictive capabilities. Future research in this area may further explore various data collection strategies and analytical techniques to develop a more comprehensive framework for selecting and utilizing proxy models effectively.
Ethical Considerations
We are committed to conducting our evaluations with the utmost standards of transparency and fairness. This commitment involves applying rigorous methodologies that ensure equitable and unbiased assessment processes.
Acknowledgements
We extend our sincere gratitude to Viktoria Schram for providing assistance in reproducing Matrix Factorization as our baseline.
Appendix A Dataset Features Calculation
In this section, we will describe in detail how the dataset features were computed.
  • Train size: The number of dataset entries used for training specified NLP task.
  • Vocab size: The number of unique tokens in the dataset after tokenization and preprocessing, where we use SentencePiece.
  • Average Sentence Length: The average number of tokens per sentence in the dataset.
  • Word Overlap:
    $\dfrac{|T_{1}\cap T_{2}|}{|T_{1}|+|T_{2}|}$
    where $T_{1}$ and $T_{2}$ correspond to the unique tokens of two datasets.
  • Type-Token Ratio (TTR): The ratio of unique tokens to the total number of tokens in the dataset. This is given by:
    $\text{TTR}=\dfrac{|V|}{\sum_{i=1}^{N}|s_{i}|}$
    where $|V|$ is the number of unique tokens, and $\sum_{i=1}^{N}|s_{i}|$ is the total number of tokens in all $N$ sentences.
  • TTR Distance: The distance between the TTRs of two datasets, calculated as:
    $D_{\text{TTR}}=\left(1-\dfrac{\text{TTR}_{1}}{\text{TTR}_{2}}\right)^{2}$
    where $\text{TTR}_{1}$ and $\text{TTR}_{2}$ are the TTR values for two datasets.
  • Jensen-Shannon Divergence (JSD): Measures the divergence between two token distributions. For two token distributions $P$ and $Q$, the JSD is calculated as:
    $\text{JSD}(P,Q)=\dfrac{1}{2}\left[\text{KL}(P\parallel M)+\text{KL}(Q\parallel M)\right]$
    where $M=\dfrac{1}{2}(P+Q)$ and KL is the Kullback-Leibler divergence:
    $\text{KL}(P\parallel Q)=\sum_{x}P(x)\log\dfrac{P(x)}{Q(x)}$
  • TF-IDF Cosine Similarity: The cosine similarity between two datasets based on their TF-IDF representations. The cosine similarity is given by:
    Cosine $\text{Cosine Similarity}(A,B)=\dfrac{A\cdot B}{\|A\|\|B\|}$
    where $A$ and $B$ are the TF-IDF vectors of the two datasets.
  • Sentence-BERT Similarity: The cosine similarity between the mean sentence embeddings of two datasets obtained from Sentence-BERT Reimers and Gurevych (2019). It is computed as:
    Cosine $\text{Cosine Similarity}(E_{A},E_{B})=\dfrac{E_{A}\cdot E_{B}}{\|E_{A}\|\|E_{B}\|}$
    where $E_{A}$ and $E_{B}$ are the mean sentence embeddings for datasets $A$ and $B$, respectively.
Appendix B Experimental Details
Datasets Languages Under Study Domain
English-centric Dataset
   FLoRes-200 Costa-jussà et al. (2022) afr, amh, arz, bak, bel, ceb, cym, deu, dik, ewe, Multi-domain
fao, fra, guj, hau, hne, hye, ibo, ind, jav, kan,
kat, kaz, khm, kin, kir, kor, lmo, ltz, mar, mri,
mya, oci, pan, plt, rus, sin, sna, snd, som, ssw,
tam, tat, tgl, tuk, wol, xho, vie, yor, zho, zul
   MT560 Gowda et al. (2021)
      Europarl deu Wiki
      DGT fra Government
      Joshua Indian Corpus tam Wiki
      Neulab TED Talk Qi et al. (2018) bel, deu, fra, hye, kat, kaz, kor, mar, mya, rus, TED
tam, vie, zho
      News Commentary deu, fra, rus, zho News
      News-test WMT22  Kocmi et al. (2022) deu, rus, zho News
      General-test WMT22  Kocmi et al. (2022) deu, fra, kaz, rus, zho News
      OPUS100 Zhang et al. (2020) afr, amh, bel, cym, deu, fra, guj, hau, hye, ibo, Multi-domain
ind, kan, kat, kaz, khm, kin, kir, kor, mar, mya,
oci, pan, rus, sin, tam, tat, tuk, vie, xho, yor,
zho, zul
      OPUS Bible afr, amh, ceb, deu, dik, ewe, fra, guj, hye, ind, Religion
      Christodouloupoulos and Steedman (2015) kan, kor, mar, mri, mya, plt, rus, sin, sna, som,
ssw, tam, tgl, vie, wol, xho, zho, zul
      OPUS OpenSubtitles afr, hye, kat, kor, sin, tgl, vie, zho Movies
      OPUS Tatoeba Tiedemann (2012) afr, amh, arz, bak, bel, ceb, fao, hau, hye, ibo, Conversational
ind, jav, kan, kat, kaz, khm, kin, kir, kor, lmo,
ltz, mar, mri, mya, pan, sna, tat, tgl, tuk, vie,
yor, zul
      OPUS Tanzil Tiedemann (2012) amh, deu, fra, hau, ind, kor, rus, snd, som, tat, Religion
zho
      OPUS Gnome Tiedemann (2012) fao, mri, som, zho Technology
      OPUS GlobalVoices Tiedemann (2012) amh, kor, mya News
      OPUS Wikipedia Health Tiedemann (2012) fra, kor, rus, vie Health
      OPUS Antibiotic Tiedemann (2012) fra Health
      OPUS Tico Tiedemann (2012) fra, rus Health
      OPUS Vaccination Tiedemann (2012) fra Health
      OPUS Ubuntu Tiedemann (2012) vie, zho Technology
      OPUS EU Bookshop Tiedemann (2012) cym Government
      OPUS SPC afr Government
      OPUS Memat xho Medicine
      OPUS XhosaNavy xho Government
      OPUS KDE4 hne Technology
      OPUS infopankki som Immigration
      OPUS TLDR zho General
      UN rus Government
      lindat khresmoi Dušek et al. (2014) fra Health
      PMIndia Haddow and Kirefu (2020) guj, kan, mar, pan, sin, tam Government
      Wiki Titles guj, kaz, tam Wiki
      WikiMatrix Schwenk et al. (2021) arz, bak, bel, ceb, fao, fra, ind, jav, kat, kaz, Wiki
kor, lmo, ltz, mar, oci, sin, tam, tat, tgl, vie,
zho
Many-to-Many Languages Dataset
   NusaTranslation Cahyawijaya et al. (2023) bew, btk, ind, jav, mad, mak, min, sun Social Media
Table 6: List of datasets under study covering 50 different languages. We only opt for 50 out of 500 languages available in the MT560 dataset, 50 out of 200 languages available in the FLoRes-200 dataset, and 8 out of 12 languages available in the NusaTranslation dataset.
B.1 Languages Under Study
We list all the languages used in the training from the MT560 Gowda et al. (2021) and NusaTranslation Cahyawijaya et al. (2023) datasets in Table 7 and Table 9, respectively. The language code follows ISO639-3 coding. All languages are also complemented by their rarity taxonomy based on Joshi et al. (2020) into two vitality classes: 0-2→low resource language (LRL), 3-4→mid resource language (MRL), and 5→high resource language (HRL). We also provide information about whether the language was part of the pre-trained M2M100 model dataset to highlight the model knowledge coverage.
B.2 Models
The details on the proxy models we use in MT experiments are as follows:
  • Transformer (100M) Vaswani et al. (2017): a standard encoder-decoder transformer-based model with 6 encoder layers and 6 decoder layers with an embedding dimension of 512. We train the model from randomly initialized parameters with the training set.
  • SMaLL-100 (330M) Mohammadshahi et al. (2022): [^2] SMaLL-100 (330M) is taken from https://github.com/alirezamshi/small100. a distilled version of the M2M100 (12B) model. We utilize the model in two ways: fine-tuned on training data and zero-shot inference.
  • M2M100 (No FT) Fan et al. (2021): [^3] M2M100 (1.2B) is taken from https://github.com/facebookresearch/fairseq/tree/main/examples/m2m_100. a pre-trained estimated model of M2M100 (1.2B) without any fine-tuning. We run the model in a zero-shot fashion.
  • NLLB (No FT) Costa-jussà et al. (2022): [^4] NLLB (1.3B) is taken from https://github.com/facebookresearch/fairseq/tree/nllb. a pre-trained estimated model of NLLB-200 Distilled (1.3B) without any fine-tuning. We run the model in a zero-shot fashion.
The details on the proxy models we use in intent classification and slot filling experiments are as follows:
  • SmolLM (135M and 360M) Allal et al. (2024): [^5] SmolLM 135M and 360M are taken from https://huggingface.co/HuggingFaceTB/SmolLM-135M and https://huggingface.co/HuggingFaceTB/SmolLM-360M respectively a series of small LMs built on Cosmo-Corpus, a meticulously curated high-quality training dataset. SmolLM models have shown promising results when compared to other models in their size categories across various benchmarks testing common sense reasoning and world knowledge.
  • BLOOMZ (560M) Muennighoff et al. (2022): a decoder model fine-tuned from BLOOM Workshop et al. (2022) on a cross-lingual task mixture (xP3) and finds the resulting models capable of cross-lingual generalization to unseen tasks and languages.
Language Language Code Family Joshi Class Vitality Seen by M2M100 Covered by COMET-22
Afrikaans afr indo-european 3 MRL
Amharic amh afro-asiatic 2 LRL
Armenian hye indo-european 1 LRL
Bashkir bak turkic 1 LRL
Belarusian bel indo-european 3 MRL
Burmese mya sino-tibetan 1 LRL
Cebuano ceb austronesian 3 MRL
Chhattisgarhi hne indo-european 0 LRL
Chinese zho sino-tibetan 5 HRL
Dinka dik nilo-saharan 1 LRL
Egyptian Arabic arz afro-asiatic 3 MRL
Ewe ewe niger-congo 1 LRL
Faroese fao indo-european 1 LRL
Georgian kat kartvelian 3 MRL
German deu indo-european 5 HRL
Gujarati guj indo-european 1 LRL
French fra indo-european 5 HRL
Hausa hau afro-asiatic 2 LRL
Igbo ibo niger-congo 1 LRL
Indonesian ind austronesian 3 MRL
Javanese jav austronesian 1 LRL
Kannada kan dravidian 1 LRL
Kazakh kaz turkic 3 MRL
Khmer khm austro-asiatic 1 LRL
Kirghiz kir turkic 1 LRL
Kinyarwanda kin niger-congo 1 LRL
Korean kor koreanic 4 MRL
Lombard lmo indo-european 1 LRL
Luxembourgish ltz indo-european 1 LRL
Malagasy plt austronesian 1 LRL
Maori mri austronesian 1 LRL
Marathi mar indo-european 2 LRL
Occitan oci indo-european 1 LRL
Punjabi pan indo-european 2 LRL
Russian rus indo-european 4 MRL
Shona sna niger-congo 1 LRL
Sindhi snd indo-european 1 LRL
Sinhala sin indo-european 0 LRL
Somali som afro-asiatic 1 LRL
Swati ssw niger-congo 1 LRL
Tagalog tgl austronesian 3 MRL
Tamil tam dravidian 3 MRL
Tatar tat turkic 1 LRL
Turkmen tuk turkic 1 LRL
Vietnamese vie austro-asiatic 4 MRL
Welsh cym indo-european 1 LRL
Wolof wol niger-congo 2 LRL
Xhosa xho niger-congo 2 LRL
Yoruba yor niger-congo 2 LRL
Zulu zul niger-congo 2 LRL
Table 7: List of languages from the English-centric dataset on MT task, including their rarity category mapping, an indication of whether they are involved in the pre-training process for M2M100, and another indication of whether they are covered by COMET-22. Note that all languages are involved in the pre-training process for NLLB.
Language Language Code Family Joshi Class Vitality
Afrikaans afr indo-european 3 MRL
Amharic amh afro-asiatic 2 LRL
Arabic ara afro-asiatic 5 HRL
Azerbaijani aze turkic 1 LRL
Bengali ben indo-european 3 MRL
Welsh cym indo-european 1 LRL
Danish dan indo-european 3 MRL
German deu indo-european 5 HRL
Greek ell indo-european 3 MRL
English eng indo-european 5 HRL
Spanish spa indo-european 5 HRL
Persian fas indo-european 4 MRL
Finnish fin uralic 4 MRL
French fra indo-european 5 HRL
Hebrew heb afro-asiatic 3 MRL
Hindi hin indo-european 4 MRL
Hungarian hun uralic 4 MRL
Armenian hye indo-european 1 LRL
Indonesian ind austronesian 3 MRL
Icelandic isl indo-european 2 LRL
Italian ita indo-european 4 MRL
Japanese jpn japonic 5 HRL
Javanese jav austronesian 1 LRL
Georgian kat kartvelian 3 MRL
Khmer khm austro-asiatic 1 LRL
Kannada kan dravidian 1 LRL
Korean kor koreanic 4 MRL
Latvian lav indo-european 3 MRL
Malayalam mal dravidian 1 LRL
Mongolian mon mongolic 1 LRL
Malay msa austronesian 3 MRL
Burmese mya sino-tibetan 1 LRL
Norwegian Bokmål nob indo-european 1 LRL
Dutch nld indo-european 4 MRL
Polish pol indo-european 4 MRL
Portuguese por indo-european 4 MRL
Romanian ron indo-european 3 MRL
Russian rus indo-european 4 MRL
Slovenian slv indo-european 3 MRL
Albanian sqi indo-european 1 LRL
Swedish swe indo-european 4 MRL
Swahili swa niger-congo 2 LRL
Tamil tam dravidian 3 MRL
Telugu tel dravidian 1 LRL
Thai tha kra-dai 3 MRL
Tagalog tgl austronesian 3 MRL
Turkish tur turkic 4 MRL
Urdu urd indo-european 3 MRL
Vietnamese vie indo-european 4 MRL
Chinese (Simplified) zho sino-tibetan 5 HRL
Chinese (Traditional) zho sino-tibetan 5 HRL
Table 8: List of languages from the MASSIVE dataset for intent classification and slot filling, including their rarity category mapping.
B.3 Hyper-parameters
LM
Each fine-tuning and evaluation for LMs is done with an NVIDIA Tesla V100 32GB GPU.
The hyper-parameters used during fine-tuning from the English-centric and Many-to-Many Languages datasets for MT task are listed in Table 10, 11, 19, and 20 for SMaLL100, M2M100, NLLB, and Transformer models, respectively.
The common hyper-parameters used during fine-tuning for intent and slot classification tasks are listed in Table 12 for SmolLM-135M, SmolLM-360M, BLOOMZ-560M, Aya23-8B, and LLaMA3-8B respectively. The learning rate used for SmolLM-135M, SmolLM-360M, and BLOOMZ-560M is $1e^{-5}$, while the learning rate used for Aya23-8B and LLaMA3-8B is $1e^{-4}$. Since we are fine-tuning using SFT, we use the “default" template for SmolLM-135M, SmolLM-360M, and BLOOMZ-560M, while we use “cohere" for Aya23-8B and “llama3" for LLaMA3-8B. Additionally, we use LoRA for both Aya23-8B and LLaMA3-8B.
Regressor
Each regressor is trained on an AMD Ryzen Threadripper 2990WX with 128 GB of RAM and 16 threads. Regressors’ hyper-parameters used are provided in Table 21, 22, 23, 24, 25, and 26 for XGBoost, Poly2/Poly3, LGBM, and MF, respectively. These hyper-parameters were obtained based on the best cross-validation RMSE score on training using 10 folds.
Language Language Code Family Joshi Class Vitality Seen by M2M100 Seen by NLLB
Indonesian ind austronesian 3 MRL
Javanese jav austronesian 1 LRL
Betawi bew creole 0 LRL
Batak btk austronesian 0 LRL
Madurese mad austronesian 0 LRL
Makassarese mak austronesian 0 LRL
Minangkabau min austronesian 0 LRL
Sundanese sun austronesian 1 LRL
Table 9: List of languages from the Many-to-Many Languages dataset on MT task along with their rarity category mapping and an indication of whether they are included in the pre-training process for each respective model.
Hyper-parameter English-centric Many-to-Many Langs.
Encoder Layers 12 12
Decoder Layers 3 3
Encoder Embed Dim 1024 1024
Decoder Embed Dim 1024 1024
Encoder FFN Embed Dim 4096 4096
Decoder FFN Embed Dim 4096 4096
Encoder Attention Heads 16 16
Decoder Attention Heads 16 16
Encoder Layerdrop 0.05 0.05
Decoder Layerdrop 0.05 0.05
Optimizer Adam Adam
Adam Eps 1e-6 1e-6
Adam Betas (0.9, 0.98) (0.9, 0.98)
Patience 6 6
Batch Size 16 16
Dropout 0.1 0.1
Attention Dropout 0.1 0.1
ReLU Dropout 0.1 0.1
Weight Decay 0.0 0.0
Label Smoothing 0.1 0.1
Clip Norm 1.0 1.0
Learning Rate 0.0001 0.0003
Max Tokens (per GPU) 1,000 1,000
Table 10: List of hyper-parameters used for SMaLL100 with English-centric and Many-to-Many Languages datasets.
Hyper-parameter English-centric Many-to-Many Langs.
Encoder Layers 24 24
Decoder Layers 24 24
Encoder Embed Dim 1024 1024
Decoder Embed Dim 1024 1024
Encoder FFN Embed Dim 8192 8192
Decoder FFN Embed Dim 8192 8192
Encoder Attention Heads 16 16
Decoder Attention Heads 16 16
Encoder Layerdrop 0.05 0.05
Decoder Layerdrop 0.05 0.05
Optimizer Adam Adam
Adam Eps 1e-6 1e-6
Adam Betas (0.9, 0.98) (0.9, 0.98)
Patience 6 6
Batch Size 32 32
Dropout 0.1 0.1
Attention Dropout 0.1 0.1
ReLU Dropout 0.1 0.1
Weight Decay 0.0 0.0
Label Smoothing 0.1 0.1
Clip Norm 0.0 0.0
Learning Rate 0.0002 0.0002
Max Tokens (per GPU) 1,792 1,792
Table 11: List of hyper-parameters used for M2M100 with English-centric and Many-to-Many Languages datasets.
Hyper-parameter Value
Cutoff Length 256
Preprocessing Num Workers 16
Train Batch Size 1
Gradient Accumulation Steps 2
Epochs 3
Learning Rate Scheduler cosine
Warmup Ratio 0.1
Eval Batch Size 1
Eval Steps 500
Table 12: List of common hyper-parameters used for fine-tuning SmolLM-135M, SmolLM-360M, BLOOMZ-560M, Aya23-8B, and LLaMA3-8B on intent and slot classification task.
B.4 Regressor Dataset Sizes
We provide the details of the regressor’s training and test set size in Table 27.
Appendix C More Detailed Results
Models Random LOLO
[…]
NLPerf Xia et al. (2020) with different regressors
XGBoost 7.69 $\pm$ 0.59 7.73 $\pm$ 0.08 9.20 $\pm$ 0.40 12.92 $\pm$ 0.54 9.16
Poly2 Khiu et al. (2024) 11.21 $\pm$ 0.49 16.23 $\pm$ 3.51 15.55 $\pm$ 0.00 43.02 $\pm$ 0.00 34.62
Poly3 Khiu et al. (2024) 11.00 $\pm$ 0.49 15.64 $\pm$ 2.59 62.29 $\pm$ 0.00 236.29 $\pm$ 0.00 66.97
LGBM Ye et al. (2021) 7.88 $\pm$ 0.65 8.15 $\pm$ 0.19 9.71 $\pm$ 0.17 12.81 $\pm$ 0.23 9.62
ProxyLM (Ours) with different proxy models
Transformer 4.68 $\pm$ 0.41 7.22 $\pm$ 0.25 6.18 $\pm$ 0.30 11.78 $\pm$ 0.50 7.31
SMaLL-100 4.07 $\pm$ 0.31 6.33 $\pm$ 0.11 4.59 $\pm$ 0.24 10.33 $\pm$ 0.41 6.04
SMaLL-100 (No FT) 5.27 $\pm$ 0.50 6.04 $\pm$ 0.32 6.28 $\pm$ 0.32 10.94 $\pm$ 0.50 6.95
Estimated Model (No FT) 5.23 $\pm$ 0.54 4.15 $\pm$ 0.24 6.18 $\pm$ 0.31 5.42 $\pm$ 0.27 5.38
Ensemble 3.21 $\pm$ 0.29 3.68 $\pm$ 0.36 3.74 $\pm$ 0.20 4.94 $\pm$ 0.29 4.01
Table 13: English-centric test results using spBLEU in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning" and the model inference is done in a zero-shot fashion. Avg represents the average of the results across the row. The reported results use XGBoost as the regressor. Ensemble denotes combining all four proxy models.
Models Random LOLO
[…]
NLPerf Xia et al. (2020) with different regressors
XGBoost 0.0717 $\pm$ 0.0018 0.0556 $\pm$ 0.0038 0.1130 $\pm$ 0.0054 0.1007 $\pm$ 0.0046 0.0825
Poly2 Khiu et al. (2024) 0.1325 $\pm$ 0.0014 0.1503 $\pm$ 0.0019 0.1258 $\pm$ 0.0000 0.1176 $\pm$ 0.0000 0.1316
Poly3 Khiu et al. (2024) 0.1325 $\pm$ 0.0014 0.1520 $\pm$ 0.0019 0.1258 $\pm$ 0.0000 0.1176 $\pm$ 0.0000 0.1320
LGBM Ye et al. (2021) 0.0752 $\pm$ 0.0025 $0$$\pm$ 0.0033 0.1147 $\pm$ 0.0000 0.1019 $\pm$ 0.0000 0.0880
ProxyLM (Ours) with different proxy models
Transformer 0.0543 $\pm$ 0.0009 0.0492 $\pm$ 0.0041 0.0837 $\pm$ 0.0042 0.0957 $\pm$ 0.0031 0.0707
SMaLL-100 0.0328 $\pm$ 0.0012 0.0431 $\pm$ 0.0033 0.0547 $\pm$ 0.0027 0.0824 $\pm$ 0.0030 0.0533
SMaLL-100 (No FT) 0.0454 $\pm$ 0.0024 0.0393 $\pm$ 0.0033 0.0778 $\pm$ 0.0030 0.0967 $\pm$ 0.0034 0.0648
Estimated Model (No FT) 0.0460 $\pm$ 0.0017 0.0348 $\pm$ 0.0016 0.0777 $\pm$ 0.0034 0.0607 $\pm$ 0.0034 0.0548
Ensemble 0.0289 $\pm$ 0.0012 0.0293 $\pm$ 0.0013 0.0454 $\pm$ 0.0019 0.0613 $\pm$ 0.0031 0.0412
Table 14: English-centric MT test results using COMET-22 in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning" and the model inference is done in a zero-shot fashion. There are no results for the Unseen setting since COMET-22 score does not cover the unseen languages to M2M100. Furthermore, all languages are covered by NLLB in English-centric dataset. Avg represents the average of the results across the row. The reported results use XGBoost as the regressor. Ensemble denotes combining all four proxy models, the detailed breakdown of this result can be seen in Section C in the Appendix.
Models Random LOLO
[…]
NLPerf Xia et al. (2020) with different regressors
XGBoost 2.45 $\pm$ 0.30 1.11 $\pm$ 0.06 7.83 $\pm$ 0.23 8.28 $\pm$ 0.31 4.94
Poly2  Khiu et al. (2024) 4.70 $\pm$ 0.40 4.68 $\pm$ 0.51 7.07 $\pm$ 0.00 7.90 $\pm$ 0.00 6.09
Poly3  Khiu et al. (2024) 4.60 $\pm$ 0.41 4.64 $\pm$ 0.49 7.26 $\pm$ 0.00 8.01 $\pm$ 0.00 6.13
LGBM  Ye et al. (2021) 2.66 $\pm$ 0.22 1.76 $\pm$ 0.29 7.91 $\pm$ 0.01 8.06 $\pm$ 0.00 5.10
MF Schram et al. (2023) 3.65 $\pm$ 0.26 2.60 $\pm$ 0.39 7.08 $\pm$ 0.23 7.14 $\pm$ 0.22 5.12
ProxyLM (Ours) with different proxy models
Transformer 2.56 $\pm$ 0.43 1.70 $\pm$ 0.20 5.65 $\pm$ 0.23 6.24 $\pm$ 0.34 4.04
SMaLL-100 2.56 $\pm$ 0.33 1.65 $\pm$ 0.44 4.85 $\pm$ 0.36 5.14 $\pm$ 0.46 3.55
SMaLL-100 (No FT) 2.44 $\pm$ 0.21 1.34 $\pm$ 0.38 6.93 $\pm$ 0.34 7.25 $\pm$ 0.37 4.49
Estimated Model (No FT) 2.38 $\pm$ 0.36 1.27 $\pm$ 0.03 5.10 $\pm$ 0.28 5.50 $\pm$ 0.26 3.56
Ensemble 2.41 $\pm$ 0.28 1.56 $\pm$ 0.35 3.73 $\pm$ 0.23 3.79 $\pm$ 0.19 2.90
Table 15: Many-to-Many Languages test results in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning". Avg represents the average of the results across the row. The reported results are experiments using XGBoost regressor. Ensemble denotes combining all four proxy models.
Models Unseen Cross-Dataset
[…]
NLPerf Xia et al. (2020) with different regressors
XGBoost 8.26 $\pm$ 0.53 12.90 $\pm$ 1.01 12.02 $\pm$ 1.02
Poly2  Khiu et al. (2024) 9.51 $\pm$ 0.00 11.02 $\pm$ 0.00 8.97 $\pm$ 0.00
Poly3  Khiu et al. (2024) 9.64 $\pm$ 0.00 11.06 $\pm$ 0.00 10.98 $\pm$ 0.00
LGBM  Ye et al. (2021) 9.56 $\pm$ 0.97 9.34 $\pm$ 0.00 10.58 $\pm$ 0.00
ProxyLM (Ours) with different proxy models
Transformer 6.71 $\pm$ 0.37 9.31 $\pm$ 0.00 9.70 $\pm$ 0.00
SMaLL-100 4.87 $\pm$ 0.34 4.52 $\pm$ 0.05 6.87 $\pm$ 0.11
SMaLL-100 (No FT) 6.22 $\pm$ 0.38 8.26 $\pm$ 0.59 9.50 $\pm$ 0.04
Estimated Model (No FT) 5.90 $\pm$ 0.40 9.95 $\pm$ 0.40 7.20 $\pm$ 0.24
Ensemble 4.48 $\pm$ 0.23 4.35 $\pm$ 0.02 5.03 $\pm$ 0.48
Table 16: Unseen and Cross-Dataset MT test results on English-centric dataset in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. The columns show the setting and estimated model. “No FT" denotes “no fine-tuning". We only show M2M100 results for the Unseen setting since NLLB covers all languages in the English-centric dataset. The reported results for the Unseen setting use XGBoost, while the Cross-Dataset experiments use LGBM. Ensemble denotes combining all four proxy models.
Models Random LOLO
Aya-23$\downarrow$ LLaMA3$\downarrow$ Aya-23$\downarrow$ LLaMA3$\downarrow$ Avg.
NLPerf Xia et al. (2020) with different regressors
XGBoost 0.0761 $\pm$ 0.0008 0.0191 $\pm$ 0.0008 0.1573 $\pm$ 0.0039 0.0581 $\pm$ 0.0021 0.0777
Poly2 Khiu et al. (2024) 0.1996 $\pm$ 0.0017 0.0979 $\pm$ 0.0033 0.2075 $\pm$ 0.0000 0.0918 $\pm$ 0.0000 0.1492
Poly3 Khiu et al. (2024) 0.1990 $\pm$ 0.0011 0.0969 $\pm$ 0.0015 0.2191 $\pm$ 0.0000 0.0925 $\pm$ 0.0000 0.1519
LGBM Ye et al. (2021) 0.0839 $\pm$ 0.0030 0.0198 $\pm$ 0.0023 0.1545 $\pm$ 0.0000 0.0558 $\pm$ 0.0006 0.0785
ProxyLM (Ours) with different proxy models
SmolLM (135M) 0.0676 $\pm$ 0.0013 0.0171 $\pm$ 0.0008 0.1273 $\pm$ 0.0033 0.0455 $\pm$ 0.0020 0.0644
SmolLM (360M) 0.0604 $\pm$ 0.0015 0.0157 $\pm$ 0.0003 0.1118 $\pm$ 0.0031 0.0441 $\pm$ 0.0017 0.0580
BLOOMZ (560M) 0.0692 $\pm$ 0.0025 0.0179 $\pm$ 0.0009 0.1283 $\pm$ 0.0038 0.0482 $\pm$ 0.0023 0.0659
Ensemble 0.0609 $\pm$ 0.0010 0.0164 $\pm$ 0.0008 0.1112 $\pm$ 0.0032 0.0442 $\pm$ 0.0019 0.0582
Table 17: Intent classification results using accuracy in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. Avg represents the average of the results across the row. The reported results use XGBoost as the regressor. Ensemble denotes combining all four proxy models.
Models Random LOLO
Aya-23$\downarrow$ LLaMA3$\downarrow$ Aya-23$\downarrow$ LLaMA3$\downarrow$ Avg.
NLPerf Xia et al. (2020) with different regressors
XGBoost 0.0693 $\pm$ 0.0009 0.0548 $\pm$ 0.0034 0.1219 $\pm$ 0.0035 0.1093 $\pm$ 0.0027 0.0888
Poly2 Khiu et al. (2024) 0.1396 $\pm$ 0.0008 0.1412 $\pm$ 0.0017 0.1418 $\pm$ 0.0000 0.1414 $\pm$ 0.0000 0.1410
Poly3 Khiu et al. (2024) 0.1393 $\pm$ 0.0008 0.1401 $\pm$ 0.0017 0.1448 $\pm$ 0.0000 0.1413 $\pm$ 0.0000 0.1414
LGBM Ye et al. (2021) 0.0692 $\pm$ 0.0017 0.0557 $\pm$ 0.0029 0.1218 $\pm$ 0.0000 0.1152 $\pm$ 0.0000 0.0905
ProxyLM (Ours) with different proxy models
SmolLM (135M) 0.0618 $\pm$ 0.0018 0.0538 $\pm$ 0.0022 0.1004 $\pm$ 0.0028 0.0953 $\pm$ 0.0024 0.0778
SmolLM (360M) 0.0562 $\pm$ 0.0011 0.0506 $\pm$ 0.0023 0.0844 $\pm$ 0.0024 0.0868 $\pm$ 0.0018 0.0695
BLOOMZ (560M) 0.0618 $\pm$ 0.0027 0.0540 $\pm$ 0.0019 0.1023 $\pm$ 0.0033 0.0995 $\pm$ 0.0028 0.0794
Ensemble 0.0561 $\pm$ 0.0016 0.0508 $\pm$ 0.0021 0.0830 $\pm$ 0.0022 0.0884 $\pm$ 0.0025 0.0696
Table 18: Slot filling results using micro-F1 score in average RMSE $\pm$ standard deviation (lower is better). Bold numbers indicate the best performance, while underlined numbers represent the second-best performance. Avg represents the average of the results across the row. The reported results use LGBM as the regressor. Ensemble denotes combining all four proxy models.
We provide detailed results for Table 1, 2, and 2 by providing the standard errors of the predictions. The mapping of vitality, Joshi class, and language family follows the classifications in Table 7 and 9. The mapping of all languages in Table 5 until 7.
Appendix D Feature Importances
We provide feature importance scores of XGBoost with ProxyLM Ensemble for the random English-centric experiment in Figure 8 and 9. Each combination consists of one most influential feature followed by others with marginal contributions to the model, each with an importance score of 0.12 or less. We observe that proxy models are always the most influential features in prediction. Other feature importances plot are also provided in Figure 11 until 13.
Appendix E License for Artifacts
We discuss the license or terms for the use of any artifacts we use in Table 28.
Hyper-parameter English-centric Many-to-Many Langs.
Encoder Layers 24 24
Decoder Layers 24 24
Encoder Embed Dim 1024 1024
Decoder Embed Dim 1024 1024
Encoder FFN Embed Dim 8192 8192
Decoder FFN Embed Dim 8192 8192
Encoder Attention Heads 16 16
Decoder Attention Heads 16 16
Encoder Layerdrop 0.05 0.05
Decoder Layerdrop 0.05 0.05
Optimizer Adam Adam
Adam Eps 1e-6 1e-6
Adam Betas (0.9, 0.98) (0.9, 0.98)
Patience 6 6
Batch Size 32 32
Dropout 0.1 0.1
Attention Dropout 0.1 0.1
ReLU Dropout 0.0 0.0
Weight Decay 0.01 0.01
Label Smoothing 0.1 0.1
Clip Norm 1.0 1.0
Learning Rate 0.00002 0.0001
Max Tokens (per GPU) 1,000 1,000
Table 19: List of hyper-parameters used for NLLB with English-centric and Many-to-Many Languages datasets.
Hyper-parameter English-centric Many-to-Many Langs.
Encoder Layers 6 6
Decoder Layers 6 6
Encoder Embed Dim 512 512
Decoder Embed Dim 512 512
Encoder FFN Embed Dim 2048 2048
Decoder FFN Embed Dim 2048 2048
Encoder Attention Heads 8 8
Decoder Attention Heads 8 8
Encoder Layerdrop 0.05 0.05
Decoder Layerdrop 0.05 0.05
Optimizer Adam Adam
Adam Eps 1e-6 1e-6
Adam Betas (0.9, 0.98) (0.9, 0.98)
Patience 6 6
Batch Size 32 32
Dropout 0.1 0.1
Attention Dropout 0.1 0.1
ReLU Dropout 0.1 0.1
Weight Decay 0.0001 0.0001
Label Smoothing 0.1 0.1
Clip Norm 0 0
Learning Rate 0.001 0.0005
Max Tokens (per GPU) 1,000 1,000
Table 20: List of hyper-parameters used for Transformer with English-centric and Many-to-Many Languages datasets.
Hyper-parameter Using spBLEU Using COMET
M2M100 NLLB M2M100 NLLB
max n_estimators 5000 5000 5000 5000
max eta 0.1 0.1 0.1 0.1
min_child_weight 5.0 4.2 3.2 1.1
max_depth 5 4 3 5
gamma 0 0 0 0
subsample 0.6 0.94 0.6 1
colsample_bytree 0.83 0.82 0.9 0.86
reg_alpha 0.2 0.32 0.11 0
reg_lambda 0.1 0.37 0.48 0.05
Table 21: List of hyper-parameters used for XGBoost Regressor on MT task with M2M100 and NLLB models trained with English-centric dataset.
Hyper-parameter M2M100 NLLB
max n_estimators 2000 2000
max eta 0.1 0.1
min_child_weight 5 2.5
max_depth 3 3
gamma 0 0
subsample 0.7 0.9
colsample_bytree 0.6 0.6
reg_alpha 0 0
reg_lambda 0.35 0.15
Table 22: List of hyper-parameters used for XGBoost Regressor on MT task with M2M100 and NLLB models trained with Many-to-Many Languages dataset.
Hyper-parameter Aya-23 LLaMA3
max n_estimators 5000 5000
max eta 0.1 0.1
min_child_weight 3.0 3.0
max_depth 3 3
gamma 0.1 0.1
subsample 0.85 0.6
colsample_bytree 1.0 0.95
reg_alpha 0.1 0.1
reg_lambda 0.2 0.5
Table 23: List of hyper-parameters used for XGBoost Regressor on intent classification and slot filling tasks with Aya-23 and LLaMA-3 models.
Hyper-parameter Value
alpha 0.1
l1_ratio 0.9
Table 24: List of hyper-parameters used for Poly2/Poly3 Regressor for all tasks.
Hyper-parameter Value
max learning_rate 0.3
max num_leaves 64
n_estimators 100
max_bin 200000
max_depth 10
min_child_weight 0.001
min_child_samples 20
min_split_gain 0.0
colsample_bytree 1.0
subsample 1.0
reg_alpha 0.1
reg_lambda 0.1
Table 25: List of hyper-parameters used for LGBM Regressor for all tasks. “Max" indicates the maximum value set for the hyper-parameter during the hyper-parameter search.
Hyper-parameter Specification
max alpha 0.01
beta_w 0.1
beta_h 0.1
beta_z 0.01
beta_s 0.01
beta_t 0.01
lr_decay 0.001
iterations 2000
Table 26: List of hyper-parameters used for MF Regressor with M2M100 and NLLB models trained with Many-to-Many Languages datasets. “Max" indicates the maximum value set for the hyper-parameter during the hyper-parameter search.
Experimental Settings Train Size Test Size
Random (English-centric) 1,367 587
Random (Many-to-Many Langs.) 156 68
Random (Intent Classification) 1,820 781
Random (Slot Filling) 1,820 781
Unseen 1,853 101
Cross-Dataset 1,954 224
Table 27: Regressor’s training and test set size on different experimental settings. The total MT experimental records for English-centric and Many-to-Many Languages datasets are 1,954 and 224, respectively. On the other hand, the total experimental records for intent classification and slot filling are both 2,601.
Datasets URL Link License
MT560 Gowda et al. (2021) https://opus.nlpl.eu/MT560 Unknown
FLoRes Costa-jussà et al. (2022) Muennighoff/flores200 CC-BY-SA 4.0
NusaTranslation Cahyawijaya et al. (2023) https://huggingface.co/datasets/indonlp/nusatranslation_mt Apache 2.0
MASSIVE FitzGerald et al. (2022) https://huggingface.co/datasets/AmazonScience/massive CC-BY 4.0
Table 28: List of datasets under study with their licenses.
Figure 5: Detailed results of XGBoost with ProxyLM Ensemble on the M2M100 model under the LOLO setting using the English-centric dataset on MT task from Table 13 per languages.
Figure 6: Detailed results of XGBoost with ProxyLM Ensemble on the LLaMA3 model under the LOLO setting on intent classification from Table 17 per languages.
Figure 7: Detailed results of LGBM with ProxyLM Ensemble on the Aya model under the LOLO setting on slot classification from Table 18 per languages.
Figure 8: Feature importance analysis of XGBoost with ProxyLM Ensemble on MT task using M2M100 model using the English-centric dataset.
Figure 9: Feature importance analysis of XGBoost with ProxyLM Ensemble on MT task using NLLB model using the English-centric dataset.
Figure 10: Feature importance analysis of XGBoost with ProxyLM Ensemble on the Aya-23 on slot filling task.
Figure 11: Feature importance analysis of XGBoost with ProxyLM Ensemble on the LLaMA3 on slot filling task.
Figure 12: Feature importance analysis of XGBoost with ProxyLM Ensemble on the Aya-23 on intent classification task.
Figure 13: Feature importance analysis of XGBoost with ProxyLM Ensemble on the LLaMA3 on intent classification task.